### v1

# The mantle-inner core gravitational mode of oscillation in a strong magnetic field regime

Abstract The mantle-inner core gravitational (MICG) mode is the free mode axial oscillation between the mantle and inner core sustained by the gravitational torque between their degree 2 order 2 density structures. Here, we investigate how the MICG mode is affected by oscillations of cylindrical surfaces in the fluid outer core in the form of Alfvén waves. The latter are triggered by oscillations of the tangent cylinder (TC) moving jointly with the inner core and propagate away from the rotation axis. We show that the MICG mode remains a distinct normal mode of oscillation of the core-mantle system only when the triggered Alfvén waves are attenuated before they traverse the width of the fluid core. For an internal magnetic field strength of a few mT, as we expect in Earth’s core, Alfvén waves can readily traverse the width of the core, and the MICG mode is absorbed into the spectrum of torsional oscillation (TO) modes. The MICG period retains a dynamical influence, acting as a point of resonance for TO modes, and marking the transition from a TO mode in which the motion of the TC (including the inner core) is weakly impacted by gravitational coupling to one in which the oscillating motion of the TC is strongly restricted. Our results imply that the observed 6-year periodic signal in the length of day cannot be interpreted as the signature of the MICG mode and must instead be caused by TO modes, or more generally, by the propagation of Alfvén waves. {summary}Non-technical summary The Earth’s mantle is not uniform but contains continental size density anomalies. The topography of the solid inner core (located at the centre of our planet) is aligned on average with these mantle density anomalies. The axial gravitational restoring torque between these mantle and inner core structures – when they are rotated out of longitudinal alignment – sustains a natural mode of oscillation termed the mantle-inner core gravitational (MICG) mode. Previous studies have suggested that the MICG mode may explain a periodic signal of 6 years observed in the changes in the length of day (LOD). Here, we show that when the magnetic field inside the fluid core is high (of the order of a few mT), the nature of the MICG changes; it is no longer a distinct, independent mode of oscillation of the core-mantle system. Instead, the MICG mode merges into the set of modes of torsional oscillations (TO), which consist in the longitudinal oscillations of nested cylindrical surfaces within the fluid core maintained by the magnetic tension between them (oscillations that also known as Alfvén waves). The inner core can still oscillate with respect to the mantle, but it is due to its entrainment as part of the TO modes. The MICG mode retains a dynamical influence, as its period determines whether the inner core oscillation for a given TO mode is impeded or not by the mantle gravitational torque. These results have important implications for the interpretation of the 6 year LOD signal, which cannot be caused by the MICG mode, but must instead be caused by the TO modes, or more generally, by the propagation of Alfvén waves in the fluid core.
1 Introduction
The mantle-inner core gravitational (MICG) mode of axial oscillation owes its existence to aspherical mass anomalies within the mantle, including those caused by the topography at surfaces of density contrast (buffett96a). The largest of these occur at spherical harmonic degree 2 order 2 (e.g. defraigne96; simmons07), warping equipotential surfaces everywhere inside the Earth (including the surface geoid) into ellipsoid shapes elongated in the equatorial plane. Within the core, surfaces of constant density coincide with these elliptically shaped equipotential surfaces in order to maintain hydrostatic equilibrium. Provided the inner core deforms viscously on a timescale shorter than the typical changes in the gravitational potential imposed by the mantle, the inner core boundary (ICB) should also match an equipotential surface. Its shape should then be elliptical in the plane of the equator (i.e. equatorially elliptical) and longitudinally aligned with the leading order mantle mass anomalies (buffett96a). An axial rotation of the inner core out of this longitudinal alignment, driven for example by magnetic stresses at the ICB (e.g. buffett00; aubert11), leads to a gravitational torque acting to restore the equilibrium and state of minimum gravitational potential energy (buffett96a). This gravitational torque can thus sustain a natural mode of axial oscillation between the inner core and the mantle, referred to as the MICG mode (mound06).
The period of the MICG mode depends on the gravitational strength factor, $\overline{\Gamma}$, capturing the amplitude of the gravitational torque for a given small longitudinal misalignment between the inner core and mantle (buffett96b). In turn, $\overline{\Gamma}$ depends on the amplitude of the degree 2 order 2 geoid at the core-mantle boundary (CMB), which itself depends on estimates of the mantle density field, including that from the topography of the CMB (buffett96a; davies14; chao17; shih21). The period of the MICG mode also depends on whether the region of the fluid core inside the axial cylinder that encloses the inner core – the so-called tangent cylinder (TC) – is entrained into co-rotation with the inner core (see eqs. (8) and (9) below).
Estimates of the period of the MICG mode can be derived by computing predictions of $\overline{\Gamma}$ on the basis of viscous mantle flow models. The latter are built from estimates of the density anomalies within the mantle inferred from seismic tomography and the subsequent viscous deformation of the CMB, the surface and other surfaces of density discontinuities that they cause (e.g. hager85; forte94; defraigne96). Based on the mantle flow models available at the time, buffett96a estimated a MICG period of approximately 2 to 3 years. The more recent survey of davies14 suggests a longer period in the range 7 to 18 years. The range of uncertainty reflects our limited knowledge of both the amplitude of the internal mantle density anomalies (primarily those of degree 2, order 2 associated with Large Low Seismic Velocity provinces in the lower mantle, e.g. garnero16; mcnamara19) and the viscosity of the lowermost mantle.
An alternate approach is to seek the signature of the MICG mode in the observed changes in the mantle rotation rate measured as changes in the length of day (LOD). Indeed, an observed quasi-periodic LOD signal with an amplitude of $\sim 0.13$ ms and a period of $\sim 5.9$ days (abarcadelrio00; holme13; chao14) has been linked with the MICG mode (mound06; chao17; ding18; duan18; shih21). The magnitude of $\overline{\Gamma}$ required to match this period provides then a direct measure of the equatorial ellipticities of the geoid at the CMB and of the ICB topography (davies14; chao17; shih21).
This simplified picture of the MICG mode neglects the dynamical influence of the fluid core in the region outside the tangent cylinder (ROTC) or assumes that it plays a passive role. At periods relevant to the MICG, axisymmetric azimuthal (zonal) flows are expected to be invariant in the direction of the rotation axis and to take the form of rigidly rotating nested cylindrical surfaces (jault08). A differential motion between adjacent cylinders shears the cylindrically radial ($B_{s}$) magnetic field lines that threads their surfaces. This creates an azimuthal restoring force between cylinders and allows for the propagation of Alfvén waves propagating in the cylindrically radial direction (braginsky70).
The normal modes of axial oscillations between the cylinders formed by the superposition of these Alfvén waves are referred to as torsional oscillations (TO, braginsky70) and their periods depend the strength of $B_{s}$ within the core. The prevalent view at the time the earliest studies on the MICG mode were published (e.g. buffett96a; buffett96b; mound06) was that the gravest mode of TO had a period of 60-80 years, which then implied a typical $B_{s}$-field strength of 0.2-0.3 mT within the core (zatman97). Under this scenario, a subdecadal MICG mode has a distinctly shorter period, so the ROTC may be considered to play a passive role even if it is engaged in the MICG to some degree by the propagation of Alfvén waves (e.g. mound03; mound06).
The general consensus on the magnetic field strength within the core has since shifted. Numerical models of the dynamo suggest that the field deep inside the core is likely stronger by a factor 10 than that at the CMB, suggesting an internal field strength of 3-4 mT or higher (e.g. christensen06; aubert09). At such a strength, Alfvén waves should propagate through the entire width of the core in only a few years and the periods of the gravest TO modes are expected to be in the subdecadal range. Indeed, gillet10 showed that the angular momentum carried by propagating Alfvén waves in the ROTC matches that required to explain the observed periodic $5.9$-yr LOD signal.
This latter result indicates that, for the 5.9 yr LOD signal, the mantle exchanges its angular momentum primarily with the ROTC. This is in contrast to the oscillation behaviour expected for the MICG mode, where the mantle exchanges its angular momentum primarily with the inner core (whether entraining the entire TC or not) and not the ROTC. It also presents an alternative explanation for the 5.9 yr LOD signal, which would then represents the entrainment of the mantle into oscillation as a result of these propagating Alfvén waves. Or, at the very least, if the MICG mode remains connected to the 5.9 yr LOD signal, the ROTC is not passive and is instead entrained into motion by the propagation of Alfvén waves.
Here, our goal is to investigate how the MICG mode is altered when a strong $B_{s}$-field inside the core allows for the propagation of fast Alfvén waves that can travel through the width of the core in a timescale shorter or similar to the MICG period. We shall show that the MICG mode is no longer an independent oscillation mode of the coupled core-mantle system but rather gets absorbed into the spectrum of TO modes, although the latter may be modified by gravitational coupling.
2 Theory
2.1 Angular momentum balance
The model of internal coupling between the mantle, fluid core and inner core that we use follows broadly that presented in several studies, including in buffett96b; buffett98; mound03; mound05b; dumberry08c; dumberry10b. We assume a spherical model of Earth’s interior comprised of an inner core (radius $r_{i}$), a fluid core (outer radius $r_{f}$) and mantle. We adopt a simplified model of flow in the fluid core restricted to its azimuthal component $v_{\phi}$ and assumed to be axisymmetric and invariant along the direction of the rotation axis. Accordingly, we use cylindrical coordinates ($s$, $\phi$, $z$), and we divide the fluid core in a set of $N_{s}$ discrete cylindrical surfaces with radius $s$ that intersect the CMB at axial position $z_{f}=\sqrt{r_{f}^{2}-s^{2}}$ in the Northern hemisphere. Cylinders inside the TC ($s<r_{i}$) intersect the ICB at $z_{i}=\sqrt{r_{i}^{2}-s^{2}}$; we set $z_{i}=0$ for cylinders outside the TC. The azimuthal velocity of each cylinder is set equal to $v_{\phi}=s\Omega_{f}$, where $\Omega_{f}\equiv\Omega_{f}(s)$ is the angular velocity.
In the absence of an external torque, the axial angular momentum equations of the mantle, inner core and fluid core are given, respectively, by
$C_{m}\frac{d}{dt}\Omega_{m}$ $=\overline{\Gamma}\alpha+\Gamma_{cmb}\,,$ (1a)
$C_{i}\frac{d}{dt}\Omega_{i}$ $=-\overline{\Gamma}\alpha+\Gamma_{icb}\,,$ (1b)
[…]
where, $C_{m}$, $C_{i}$ are the axial moments of inertia of the mantle and inner core, $\Omega_{m}$, $\Omega_{i}$ their rotation rates, and
$c_{f}=4\pi\rho s^{3}(z_{f}-z_{i})\,,$ (2)
is the axial moment of inertia density of cylinders within the fluid core, where $\rho$ is the density (assumed uniform). $\Gamma_{cmb}$ and $\Gamma_{icb}$ are the torques from all surface forces acting on the mantle at the CMB and on the inner core at the ICB, respectively. $\overline{\Gamma}$ is the strength of gravitational torque between the mantle and inner core and $\alpha$ is the (assumed small) longitudinal angle of misalignment between the degree 2 order 2 mantle density field and the long equatorial axis of the ICB topography. $\alpha$ depends on the differential rotation between the inner core and the mantle and on the viscous relaxation time ${\tau}$ for the ICB topography to return to its equilibrium alignment with the imposed gravitational potential from the mantle; its evolution is given by
$\frac{d\alpha}{dt}=\Omega_{i}-\Omega_{m}-\frac{\alpha}{\tau}\,.$ (3)
We assume that $\Gamma_{icb}$ and $\Gamma_{cmb}$ are both dominated by electromagnetic (EM) forces, and expressed as (e.g. buffett98)
[…]
where the coupling parameters ${\cal F}_{m}(s)$ and ${\cal F}_{i}(s)$ depend on the morphology of the magnetic field and the electrical conductivity of the solid side of the boundary. These are given by
${\cal F}_{m}(s)$ $=2\pi s^{3}G_{m}\left(\frac{z_{f}}{r_{f}}\left(B_{m}^{d}\right)^{2}+\frac{r_{f}}{z_{f}}\langle B_{r,m}\rangle^{2}\right)\,,$ (5a)
${\cal F}_{i}(s)$ $=4\pi s^{3}G_{i}\frac{z_{i}}{r_{i}}\left(B_{i}^{d}\right)^{2}\,.$ (5b)
where $B_{m}^{d}$ and $B_{i}^{d}$ are the polar amplitudes of the axial dipole at the CMB and ICB, respectively, $\langle B_{r,m}\rangle$ is the root mean squared (rms) strengths of the total radial field (including the axial dipole) at the CMB, $G_{m}$ is the conductance of the lowermost mantle and $G_{i}$ the conductance factor at the ICB given by
$G_{i}=\frac{1}{4}[1+i\,\mbox{sgn}(\omega)]\,\sigma\,\delta\,,$ (6)
where $\sigma$ is the conductivity of the fluid core (assumed equal to that of the inner core), $\delta=\sqrt{2/|\omega|\mu\sigma}$ is the magnetic skin depth at oscillation frequency $\omega$, with $\mu$ the magnetic permeability of free space, and $\mbox{sgn}(\omega)=\omega/|\omega|$. The form of ${\cal F}_{m}(s)$ was shown in dumberry08c to be a good match with the EM coupling resulting from a realistic field model at the CMB comprised of a large number of spherical harmonic terms. The field at the ICB is undoubtedly more complex than a simple axial dipole, although the latter is expected to be its dominant contribution, and it is sufficient to capture the first order effect of EM coupling at the ICB.
2.2 The MICG mode
To illustrate how the MICG mode emerges from the above system, let us assume a perfectly rigid inner core ($\tau=\infty$), and no coupling at the CMB (${\cal F}_{m}(s)=0$). Let us further assume that the fluid inside the TC rotates as a single rigid body with angular velocity $\Omega_{f}^{tc}$. Taking the time-derivatives of eqs. (1a), (1b) and (1c), and using eq. (3), gives
$C_{m}\frac{d^{2}\Omega_{m}}{dt^{2}}$ $=\bar{\Gamma}(\Omega_{i}-\Omega_{m})\,,$ (7a)
$C_{i}\frac{d^{2}\Omega_{i}}{dt^{2}}$ $=-\bar{\Gamma}(\Omega_{i}-\Omega_{m})+r_{i}\,\overline{\cal F}_{i}\frac{d}{dt}(\Omega_{f}^{tc}-\Omega_{i})\,,$ (7b)
$C_{f}^{tc}\frac{d^{2}\Omega_{f}^{tc}}{dt^{2}}$ $=-r_{i}\,\overline{\cal F}_{i}\frac{d}{dt}(\Omega_{f}^{tc}-\Omega_{i})\,,$ (7c)
where $C_{f}^{tc}$ is the moment of inertia of the fluid inside the TC and ¯$\overline{\cal F}_{i}$ is the integrated average of ${\cal F}_{i}(s)$ over the ICB. This forms a simple system of axial oscillations between $\Omega_{m}$, $\Omega_{i}$ and $\Omega_{f}^{tc}$, with a restoring gravitational torque between the inner core and mantle and a restoring (and dissipating) EM torque between the inner core and the fluid inside the TC. Note that the fluid in the ROTC is not involved and assumed to remain at rest. The natural modes of oscillations of this system include the MICG mode, whose frequency depends on the strength of $\overline{\Gamma}$ and ¯$\overline{\cal F}_{i}$.
Let us consider two end-member scenarios for ¯$\overline{\cal F}_{i}$. For no EM coupling at the ICB (¯$\overline{\cal F}_{i}=0$), the fluid inside the TC is decoupled and not involved in the MICG mode. The frequency and eigenvector of the MICG mode are (dumberry10b),
$\omega_{micg}=\sqrt{\frac{\bar{\Gamma}(C_{m}+C_{i})}{C_{m}C_{i}}}\,,\hskip 28.45274pt\left[\Omega_{m},\Omega_{i}\right]=\left[1,-\frac{C_{m}}{C_{i}}\right]\,.$ (8)
The oscillation (and angular momentum exchange) is between the mantle and inner core, with no involvement of any parts of the fluid core. As an example, taking $\overline{\Gamma}=3\times 10^{20}$ N m, and $C_{m}$ and $C_{i}$ given Table 1, this gives a MICG period of 2.8 years. If instead ¯$\overline{\cal F}_{i}$ is very large, the differential rotation between $\Omega_{f}^{tc}$ and $\Omega_{i}$ must remain very small (over the period of the MICG mode), and $\Omega_{f}^{tc}$ and $\Omega_{i}$ are locked into a common rotation. Substituting the EM coupling term of eq. (7c) into eq. (7b), and assuming $\Omega_{f}^{tc}\approx\Omega_{i}$, the frequency and eigenvector of the MICG mode are now
[…]
where $C_{tc}=C_{i}+C_{f}^{tc}$. Even though the gravitational torque remains between the mantle and inner core, the oscillation in this case is between the mantle and the whole of the TC (including the inner core). The larger moment of inertia of the whole of the TC, compared to the inner core alone, reduces the frequency of the MICG and lengthens its period. Using $\overline{\Gamma}=3\times 10^{20}$ N m as above, the period is now 5.9 years, approximately twice as long. If ¯$\overline{\cal F}_{i}$ is somewhere between these two extremes, with an amplitude such that the time it takes for $\Omega_{i}$ and $\Omega_{f}^{tc}$ to be brought back into co-rotation is of the order a few years, then the MICG mode is more complex, with each of $\Omega_{m}$, $\Omega_{i}$ and $\Omega_{f}^{tc}$ oscillating at different phases.
At periods of a few years, and assuming a radial magnetic field strength at the ICB of the order of 1 mT or larger, EM coupling is sufficiently strong to prevent any large differential rotation at the ICB (gubbins81). A rotation of the inner core will tend to entrain the fluid in contact with the ICB into co-rotation. This change in zonal flow near the ICB is communicated by the (strong) Coriolis force to the whole of the fluid column above and below the inner core, forcing it to rotate rigidly with it in accordance with the Proudman-Taylor constraint (jault08). Hence, on dynamical grounds, and for the expected strength of the magnetic field at the ICB, we expect that the MICG mode involves the whole of the TC, not just the inner core.
This simple presentation of the MICG mode is idealized in several ways. First, although ¯$\overline{\cal F}_{i}$ is most likely strong enough to lock the fluid cylinders inside the TC to the rotational motion of the inner core, some degree of differential rotation can occur near the outer edge of the TC (e.g. buffett96b; buffett98). Second, we have neglected the global elastic deformations that occur in response to the rotational displacement of the inner core out of its equilibrium longitudinal alignment. Taking these into account would lengthen the MICG period, though not by a large fraction. Third, if the inner core relaxes viscously over a relatively short time (approaching the MICG period), not only this can also lengthen the MICG period, it can lead to a significant attenuation of the mode. A fourth way this discussion of the MICG mode is idealized is that it assumes a complete decoupling of the fluid in the ROTC. Indeed, assessing the role of the ROTC on the MICG mode is the motivation for our study.
2.3 Torsional oscillation modes
The dynamical equation that governs the evolution of $\Omega_{f}$ is obtained by integrating the azimuthal component of the momentum equation over a cylindrical surface in the fluid core. For this specific geometry, the Coriolis acceleration term vanishes (e.g. taylor63; braginsky70), giving
$c_{f}\frac{\partial\Omega_{f}}{\partial t}=\frac{1}{\rho\mu}\frac{\partial}{\partial s}\left(c_{f}\{B_{s}\}\frac{b_{\phi}}{s}\right)-{\cal F}_{m}\left[\Omega_{f}-\Omega_{m}\right]-{\cal F}_{i}\left[\Omega_{f}-\Omega_{i}\right]+\nu\frac{\partial}{\partial s}\left(c_{f}\frac{\partial\Omega_{f}}{\partial s}\right)\,,$ (10)
where $\nu$ is the kinematic viscosity, $\{B_{s}\}$ is the rms strength of the $s$-component of the background magnetic field averaged over a cylindrical surface, and $b_{\phi}$ is the azimuthal magnetic field perturbation induced by the differential motion of the cylinders. The latter obeys the induction equation
$\frac{\partial b_{\phi}}{\partial t}=s\{B_{s}\}\frac{\partial\Omega_{f}}{\partial s}+\eta\left(\frac{\partial^{2}b_{\phi}}{\partial s^{2}}+\frac{1}{s}\frac{\partial b_{\phi}}{\partial s}-\frac{b_{\phi}}{s^{2}}\right)\,,$ (11)
where $\eta=1/\mu\sigma$ is the magnetic diffusivity.
When both viscous and magnetic diffusion are neglected, inserting eq. (11) into eq. (10), we retrieve the characteristic equation for Alfvén waves subject to EM friction at the top and bottom of the cylinders, (e.g. braginsky70; buffett98)
$c_{f}\frac{\partial^{2}\Omega_{f}}{\partial t^{2}}=\frac{1}{\rho\mu}\frac{\partial}{\partial s}\left(c_{f}\{B_{s}^{2}\}\frac{\partial\Omega_{f}}{\partial s}\right)-{\cal F}_{m}\frac{\partial}{\partial t}\left[\Omega_{f}-\Omega_{m}\right]-{\cal F}_{i}\frac{\partial}{\partial t}\left[\Omega_{f}-\Omega_{i}\right]\,,$ (12)
where we have assumed $\{B_{s}\}^{2}\approx\{B_{s}^{2}\}$. The normal modes of oscillations of $\Omega_{f}$ that obey eq. (12) are the torsional oscillations (TO) modes (e.g. braginsky70). In their simplest form, when assuming ${\cal F}_{m}={\cal F}_{i}=0$, these modes describe the set of azimuthal oscillations of cylinders maintained by the magnetic tension between them. The terms that contain ${\cal F}_{m}$ and ${\cal F}_{i}$ capture the coupling of the fluid cylinders with the mantle (at the CMB) and the inner core (at the ICB). Keeping these terms not only provides a way by which TO modes can exchange angular momentum with the mantle and inner core, it also influences the structure of the modes.
In all previous studies on TO modes coupled with the mantle and inner core, eq. (12) was assumed for the dynamics of $\Omega_{f}$ (buffett96b; buffett98; mound03; mound05b; dumberry08c; dumberry10b). Here, we retain both magnetic and viscous diffusion and solve for both $\Omega_{f}$ and $b_{\phi}$ using the coupled eqs. (10) and (11). The motivation for retaining magnetic diffusion is grounded in the results of luo22a who showed that it reduces the quality factor of TO modes to values below 10. Although we expect $\nu\ll\eta$ in the fluid core, the motivation for retaining viscous diffusion in (10) is for numerical convenience, as it makes it easier to prescribe appropriate boundary conditions for $\Omega_{f}$. For all our results, we have used $\nu=\eta/100$, so as to ensure that magnetic diffusion dominates over viscous diffusion.
3 Numerical implementation
Eqs. (10) and (11) are discretized on a set of grid points in the $s$-direction, with derivatives in $s$ approximated by second-order finite differences. The points $s=r_{i}=s_{tc}$ and $s=r_{f}$ are singular in eq. (10) because $dc_{f}/ds$ diverges at these locations. Likewise, the point $s=0$ is singular in eq. (11). To avoid these numerical difficulties, we set the innermost and outermost grid points at $s_{o}=0.01$ and $s_{e}=0.999$, respectively, and modify the definition of $z_{i}$ in the vicinity of $s=s_{tc}$ to
$\tilde{z}_{i}={z_{i}}-{z_{i}}\left[1-\left(\frac{s-s_{tc}}{\zeta}\right)\right]\exp\left(\frac{s-s_{tc}}{\zeta}\right)\,,$ (13)
where $\zeta\ll 1$ is an adjustable parameter. With this new definition, $d\tilde{z}_{i}/ds=0$ at $s=s_{tc}$, removing the discontinuity at the TC. The smaller $\zeta$ is, the more confined the modification is to the region close to the TC. For all our results, we used $\zeta=0.001$.
Equations (10) and (11) each require two boundary conditions, one each at $s_{o}$ and $s_{e}$. At $s=s_{o}$, we impose the regularity conditions $d\Omega_{f}/ds=0$ and $b_{\phi}=0$. At $s=s_{e}$, we set $d\Omega_{f}/ds=0$ which implies that there is no viscous torque between the fluid core and mantle and minimizes the influence of viscosity on our solutions. The condition on $b_{\phi}$ at $s_{e}$ is set by the magnetic field perturbation resulting from EM coupling (buffett98)
$b_{\phi}(s_{e})=-B_{s_{e}}\mu G_{m}s_{e}\left(\Omega_{f}(s_{e})-\Omega_{m}\right)\,,$ (14)
where $B_{s_{e}}$ is the rms amplitude of the radial magnetic field at the equator of the CMB.
With these boundary conditions, integrating eq. (10) between $s_{o}$ and $s_{e}$ yields
[…]
where $\hat{\Gamma}_{cmb}$ and $\hat{\Gamma}_{icb}$ are modified expressions for the torque on the mantle and inner core given by
[…]
The same modified expressions are used for the equations of the mantle (1a) and inner core (1b) to ensure that angular momentum is conserved.
We specify the location of the discrete $s$-grid points on two Chebyshev grids: one between $s_{o}$ and $s_{tc}$ and a second between $s_{tc}$ and $s_{e}$. The crowding of points near $s_{o}$, $s_{tc}$ and $s_{e}$ helps to resolve the thin viscous and magnetic boundary layers at $s_{o}$ and $s_{e}$ and to handle the rapidly changing $c_{f}$ in the region near the TC. We use 350 points inside the TC and 650 in the ROTC, for a total of $N_{s}$=1000 grid points, which proved to be sufficient to resolve the structure of the TO modes, and to ensure that angular momentum is conserved to within approximately 1 part in $10^{4}$.
The set of eqs. (1a), (1b), (3) and the discretized set of eqs. (10) and (11) on each of the $N_{s}$ grid points form a homogeneous system of equations for the set of $n=2N_{s}+3$ variables $\Omega_{m}$, $\Omega_{i}$, $\alpha$, $\Omega_{f}(s_{k})$ and $b_{\phi}(s_{k})$, where $k=1,2...N_{s}$. We assume that each variable has a time-dependency proportional to $\exp(-i\omega t)$, where $\omega$ is a complex frequency. The system of equations can be written in matrix form as an eigenvalue problem with eigenvalues $\omega$ and eigenvectors ${\bf x}$,
$\omega{\mathsf{B}}\cdot{\bf x}={\mathsf{A}}\cdot{\bf x}\,,$ (17)
where ${\mathsf{A}}$ and ${\mathsf{B}}$ are $n\times n$ matrices. The resulting set of modes separate into a set of $i\omega$ with negative real part (decay) and a complex conjugate imaginary part (frequency); we only solve for the positive frequencies $\omega$. Note that ${\cal F}_{i}$ in eq. (5b) has a weak dependence on $\omega$ through the magnetic skin depth $\delta$. However, for a radial magnetic field strength at the ICB of 1 mT and above, the fluid cylinders inside the TC become quasi-locked to the inner core, and this frequency dependence only has a very small effect on the normal modes. To simplify, we use a nominal frequency of $62\pi/$ 6 yr-1 for calculating $\delta$, such that ${\cal F}_{i}$ is no longer a function of $\omega$, allowing us to write the system in the simple form of eq. (17). The same strategy has been employed in previous studies (e.g. buffett98; buffett05).
Solutions depend on the choice of the profile of $\{B_{s}\}$ inside the core. We use a simple prescription given by
$\{B_{s}\}=\overline{B}_{s}\sin\left(\frac{\pi(s-s_{o})}{s_{e}-s_{o}}\right)+\left(\frac{s-s_{o}}{s_{e}-s_{o}}\right)\frac{\overline{B}_{s}}{6}\,.$ (18)
The $\sin$ term produces a maximum field of $\overline{B}_{s}$ in the bulk of the region outside the TC, similar to the $B_{s}$ profile inverted by gillet10. The linear term ensures that the radial field at the equator does not vanish, and is set such that $B_{s_{e}}=\overline{B}_{s}/6$.
The numerical values that we adopt for all other parameters are given Table 1. The dipole field amplitude $B_{m}^{d}=0.319$ mT and the total rms strength of the radial field $\langle B_{r,m}\rangle=0.391$ mT at the CMB were obtained by taking the time-average of the gufm1 model of jackson00 over the time period 1950–1990 (the rms of the radial part of the dipole field alone is 0.226 mT).
Table 1: Parameters used in calculations
Parameter value
radius of the core $r_{f}=3.480\times 10^{6}$ m
radius of the inner core $r_{i}=1.221\times 10^{6}$ m
moment of inertia, mantle $C_{m}=7.13\times 10^{37}$ kg m2
moment of inertia, inner core $C_{i}=5.87\times 10^{34}$ kg m2
moment of inertia, fluid core $C_{f}=9.42\times 10^{36}$ kg m2 (*)
moment of inertia, TC $C_{tc}=2.65\times 10^{35}$ kg m2 (*)
uniform density, fluid core $\rho=1.1\times 10^{4}$ kg m-3
conductivity, core $\sigma=10^{6}$ S m-1
magnetic diffusivity, core $\eta=\frac{1}{\mu\sigma}=0.796$ m2 s-1
kinematic viscosity, fluid core $\nu=\frac{\eta}{100}=0.00796$ m2 s-1
rms radial magnetic field, CMB $\langle B_{r,m}\rangle=0.319$ mT
axial dipole, CMB $B_{m}^{d}=0.391$ mT
axial dipole, ICB $B_{i}^{d}=3.0$ mT
(*) Based on a uniform fluid core density of $\rho=1.1\times 10^{4}$ kg m-3
4 Results
Let us first illustrate a case where the modes of axial oscillations between the mantle, inner core and fluid core include a distinct MICG mode. We set $\overline{\Gamma}=3\times 10^{20}$ N m, which gives an expected theoretical MICG frequency of $\omega_{micg}=1.065$ yr-1 (period of 5.902 years) based on eq. (9), and we set the mantle conductance at $G_{m}=10^{8}$ S. We set $\tau=100$ years to ensure that the MICG mode is only weakly damped by the viscous relaxation of the inner core. A key feature of the MICG mode is a weak involvement of the ROTC and this can be achieved by choosing a sufficiently weak $B_{s}$-field amplitude across the TC; we take $\overline{B}_{s}=0.1$ mT. We then compute the set of normal modes of oscillations for this setup, which we refer to as case $a$.
Figure 1a shows the frequency and $Q$ of a subset of the resulting modes (light blue circles) that have frequencies in the vicinity of $\omega_{micg}$. For this choice of $\overline{B}_{s}$, the fundamental mode (first harmonic) of TO has a period of $\sim 180$ years; the sequence of modes that have a $Q$ between 3 and 4 on Figure 1a correspond to higher harmonics of TO. In addition to these, a distinct mode (identified by a purple outline) with a higher $Q$ of approximately 7, separated from the TO branch, is also present: this is the MICG mode. Its frequency matches well the expected theoretical prediction under the assumption that the whole of the TC co-rotates with the inner core. Figure 2a shows the profile of $\Omega_{f}$ for this MICG mode, at the phase when $\Omega_{i}$ is maximum in the prograde direction. Inside the TC, $\Omega_{f}$ is ocked with $\Omega_{i}$ except near the edge of the TC, as found in previous studies (buffett96a; buffett98) when the radial field strength at the ICB is above a few mT. The departure of $\Omega_{f}$ from $\Omega_{i}$ near the edge of the TC implies a slight reduction in the effective moment of inertia of the TC taking part in the MICG mode and explains why its frequency is slightly higher than its theoretical prediction (see Figure 1a). Also visible in Figure 2a are the Alfvén waves generated from the TC outward. With $\overline{B}_{s}=0.1$ mT, their wave speed is sufficiently slow that they get attenuated before they reach the equator of the CMB ($s=r_{f}$).
For comparison, we also show on Figure 1a the modes (black squares) that result when there is no gravitational coupling ($\overline{\Gamma}=0$). In this case, the MICG mode is absent from the spectrum of modes; only TO modes are present, and their frequency and $Q$ coincide almost identically with those found with $\overline{\Gamma}\neq 0$. These results show that when $B_{s}$ within the core is sufficiently weak, the MICG mode emerges as an independent mode of the system, along with a set of TO modes, and the latter are only weakly influenced by gravitational coupling.
Let us now increase $B_{s}$ while keeping $\overline{\Gamma}=3\times 10^{20}$ N m and $G_{m}=10^{8}$ S. We compute the modes for $\overline{B}_{s}=0.15$ mT (case $b$) and $\overline{B}_{s}=0.3$ mT (case $c$). The results are shown on Figure 1 (panels b and c, respectively), where we again show the comparison with the modes computed with $\overline{\Gamma}=0$. As $\overline{B}_{s}$ is increased, the MICG mode is no longer a distinct, independent mode of the system. Instead, it merges into the spectrum of TO modes, although the frequency and $Q$ of these TO modes is altered by gravitational coupling. To understand why, Figures 2b and 2c show the profiles of $\Omega_{f}$ for the modes close to the MICG frequency identified by a purple outline in panels $b$ and $c$ of Figure 1. With the larger $B_{s}$-field, Alfvén waves generated at the TC travel faster and can propagate further before they get attenuated. While for case $b$ Alfvén waves are just barely able to make it through the entire width of the ROTC, they easily reach $s=r_{f}$ in case $c$. Consequently, the structure of Alfvén waves within the whole of the ROTC must be compatible with their end-condition at $s=r_{f}$ and they thus have a feedback on the oscillating motion of the TC. In other words, the oscillating TC is now taking part in TO modes, although these are modified by gravitational coupling.
Figure 1: Frequency and $Q$ of a subset of the normal modes obtained with (a) $\overline{B}_{s}=0.1$, (b) $\overline{B}_{s}=0.15$, and (c) $\overline{B}_{s}=0.3$ and different choices of coupling parameters (Pale blue circles: $\overline{\Gamma}=3\times 10^{20}$ N m, $G_{m}=10^{8}$ S; black squares: $\overline{\Gamma}=0$ N m, $G_{m}=10^{8}$; red diamonds: $\overline{\Gamma}=3\times 10^{20}$ N m, $G_{m}=10^{9}$ S). In each panel, the predicted $\omega_{micg}$ from Eq. (9) is indicated by the vertical orange line. The angular velocity structure of the modes identified by a purple outline is shown in Figure 2.
The attenuation of Alfvén waves can be achieved by reducing their speed, as illustrated by case $a$, or by increasing their attenuation rate from EM coupling at the CMB. To illustrate this, we carried out a set of experiments where we have kept $\overline{\Gamma}=3\times 10^{20}$ N m, the same three values of $\overline{B}_{s}$ as for cases $a$, $b$ and $c$ above, but used an increased mantle conductance of $G_{m}=10^{9}$ S. The modes that result from these experiments are shown on Figure 1 as red diamonds. The $Q$ of all modes is reduced because of the enhanced attenuation caused by the stronger EM coupling at the CMB. In each case though, a distinct MICG mode remains present in the mode spectrum, even with $\overline{B}_{s}=0.3$ mT.
The above analysis allows us to refine the criteria that determines whether the MICG remains an independent mode of the system. The condition is whether Alfvén waves launched by the oscillating TC get attenuated before they can reach $s=r_{f}$. When Alfvén waves are attenuated, they do not have a feedback on the motion of the TC, and their oscillating nature imply a weak net angular momentum change of the ROTC. In this sense, the ROTC plays a passive role, and the MICG mode, for which the angular momentum exchange remains dominantly between the mantle and the TC via gravitational coupling, is a possible independent mode of oscillation. In contrast, if Alfvén waves launched at the TC can travel through the entire width of the ROTC, the motion of the TC must be compatible with the allowed structure of TO modes.
Figure 2: The angular velocity of the inner core $\Omega_{i}$ (thick orange) and of the fluid core $\Omega_{f}$ (blue) as a function of cylindrical radius ($s/r_{f}$) for the modes identified by a purple outline on the respective panels of Figure 1. In each panel, the phase is chosen when $\Omega_{i}$ is maximum in the prograde direction and the amplitude is normalized such that $\max|\Omega_{f}|=1$.
Since we expect the $B_{s}$-field strength in the core to be of the order of a few mT, Alfvén waves should propagate through the ROTC in a short timescale of a couple of years, more rapidly than their timescale of attenuation by EM coupling at the CMB (which should be of the order of a few decades e.g. dumberry08c; gillet10). The MICG mode should then not exist as a distinct mode of oscillation; rather, we expect it to be absorbed into the spectrum of TO modes, although the TO modes are influenced by gravitational coupling.
To further explore how gravitational coupling affects TO modes, let us restore $G_{m}=10^{8}$ S and now set $\overline{B}_{s}=3$ mT. With these choices, the MICG mode is not an independent mode and the first TO harmonic has a period of approximately 6 years. Figure 3 shows how the period (panel a) and $Q$ (panel b) of the first four TO harmonics are altered as a function of $\overline{\Gamma}$. Figure 4 shows how the structure of the first three TO harmonics differ in a weak ($\overline{\Gamma}=10^{19}$ N m) versus a strong ($\overline{\Gamma}=10^{22}$ N m) gravitational coupling regime. The oscillation of the TC – entrained by the TO modes – is only weakly resisted by gravitational coupling when the latter is weak. In contrast, when gravitational coupling is strong, the oscillation of the TC is severely restricted. Using standing waves on a string as an analogy, a strong gravitational coupling changes the boundary condition on $\Omega_{f}$ from a free-end at $s\rightarrow 0$ to one more akin to a fixed-end at $s<s_{tc}$. Because the latter implies a shorter mode wavelength, the period of the modes is shorter, as seen in Figure 3a.
Figure 3a also shows that the point of transition of every TO mode between the weak and strong gravitational coupling regime is set by the period of the MICG mode. In other words, the TC remains gravitationally quasi-locked to the mantle for a given TO mode if its period is much longer than the MICG period. Hence, even though the MICG mode is no longer an individual mode of the system, its period retains an influence in the characteristics of the resulting TO modes. When the period of a TO mode approaches the MICG period, a resonance effect occurs, leading to a higher $Q$ (Figure 3b). The degree of resonant amplification depends on the attenuation of the TO modes caused by EM damping at the CMB but also by the viscous relaxation time of the inner core $\tau$. When the latter is of the same order or shorter than the period of a given TO mode, $Q$ is reduced. This is shown in Figure 3b, where we have computed $Q$ for $\tau=100$, $10$ and $3$ years.
Figure 3: The period (a) and $Q$ (b) of the first four TO harmonics (coloured lines, legend) as a function of $\overline{\Gamma}$. The black line in (a) is the period of the MICG mode predicted from Eq. (9). In (b), results are shown for $\tau=100$ years (solid lines), $\tau=10$ years (thin solid lines) and $\tau=3$ years (dashed lines).
Figure 4: The angular velocity as a function of cylindrical radius ($s/r_{f}$) of the (a) first, (b) second, and (c) third TO harmonics, for $\overline{\Gamma}=10^{19}$ N m (blue lines for $\Omega_{f}$, thick light blue lines for $\Omega_{i}$) and $\overline{\Gamma}=10^{22}$ N m (red lines for $\Omega_{f}$, thick orange lines for $\Omega_{i}$). In each case, the phase is chosen when $\Omega_{i}$ is maximum in the prograde direction and the amplitude is normalized such that $\max|\Omega_{f}|=1$.
5 Discussion and conclusions
As our results show, when the magnetic field strength in the bulk of the outer core is of the order of a few mT as predicted from dynamo models (christensen06; aubert09), the TC is strongly coupled by EM stress to the fluid cylinders in the ROTC. Motion in the latter take the form of Alfvén waves. If Alfvén waves launched at the TC can traverse through the entire width of the ROTC before being attenuated, the MICG mode is no longer a distinct independent mode of the coupled Earth system. Instead, it gets absorbed into the spectrum of normal modes of TO. The MICG period retains a meaning, acting as a point of resonance for TO modes, and marking the transition from a TO mode in which the oscillating motion of the TC is weakly impeded by gravitational coupling to one where the TC remains quasi-locked to the mantle.
Though presented differently, our results are in agreement with those of mound03. In their study, which uses a broadly similar model as ours, they do not compute the modes of oscillations directly. Rather, they impose a forcing on the inner core, and monitor the response of the mantle angular velocity; frequencies at which the motion of the latter is amplified mark the location of normal modes. With a uniform $B_{s}$-field of 0.4 mT (similar to our case $c$), and with no EM coupling at the CMB, they find a set of resonances which are amplified in the broad vicinity of the MICG frequency associated with TO modes (see their Figure 4). This is in agreement with our results: we also find that TO modes have a higher $Q$ when their periods is in the vicinity of the MICG period. mound03 interpreted these results to imply that the MICG mode is complicated by its interaction with TO modes. By directly computing the modes of the coupled system, we have refined this interpretation; the MICG mode is no longer a distinct independent mode, it gets integrated into the spectrum of TO modes when Alfvén waves are weakly attenuated.
With a $B_{s}$ strength in the core expected to be of the order of a few mT, Alfvén waves can traverse the core in only a few years (gillet10). For the MICG mode to remain an independent mode of the coupled Earth system, the attenuation time of Alfvén waves by coupling at the CMB (EM or otherwise) would then need to be of the order of one year or less. As an indication, such a fast attenuation would require a mantle conductance larger than $10^{9}$ S, so large in fact that the gravest TO modes and the MICG mode would all have a $Q$ below 1.
This conclusion is important regarding the interpretation of the 5.9 year oscillation in the LOD, which includes a clear contribution from core (e.g. pfeffer23; cazenave25). A number of studies have suggested that this signal may represent the signature of the MICG mode (mound06; chao17; ding18; duan18; shih21). Our results show that this interpretation, which involves the ROTC playing a passive role, cannot be correct when the magnetic field strength inside the core is a few mT. Consequently, a measure of the gravitational coupling constant $\overline{\Gamma}$ cannot be inferred on the basis of the 5.9 year LOD signal, and in turn, it cannot be used as an indirect constraint on the amplitude of the degree 2 order 2 mantle mass anomalies. Interannual LOD changes involving the core must then be the manifestation of either TO modes, or any other free or driven core flow motion. This is consistent with the results of gillet10 who showed that, for the 5.9 year LOD signal, the mantle exchanges its angular momentum with the fluid core, and dominantly the ROTC given that latter has a much larger moment of inertia than the TC.
The most recent estimates of the gravitational coupling constant $\overline{\Gamma}$ based on mantle flow models suggest a range of $0.3-2\times 10^{20}$ N m (davies14), corresponding to a range of MICG periods of 7 to 19 years. Assuming a $B_{s}$-field of approximately 3 mT or larger, all TO modes should then have a period shorter than the MICG period. This implies that TO modes should be weakly influenced by gravitational coupling in the sense that oscillations of the TC region should not be severely restricted by the mantle. Furthermore, the period of the TO modes, especially those of the first few harmonics, may be sufficiently close to the MICG period that they may be resonantly amplified. We thus expect that, at subdecadal periods, the TC (including the inner core) can undergo oscillating motion with respect to the mantle at a similar angular velocity amplitude as cylinders outside the TC, whether driven by TO modes or by any other core flow. At a period of 6 years, and based on the observed fluctuations of the magnetic field, zonal flow fluctuations inside the TC have an amplitude of approximately $v_{\phi}\sim 0.4$ km yr-1 (gillet10). Using this as a representative measure, this corresponds to longitudinal angular fluctuations of the inner core of approximately $0.02^{\circ}$ (lecomte23).
Our results also have important implications for the interpretation of the subdecadal fluctuations in inner core rotations recently inferred by seismic observations, notably in the study of wang22 which suggests a super-rotation of the inner core of $\sim 0.29^{\circ}$ between 1971 and 1974 and a sub-rotation of $\sim 0.1^{\circ}$ between 1969 and 1971. wang22 further showed that the amplitude and phase of these inferred inner core rotation changes ($\Omega_{i}$) are consistent with those reconstructed from the 5.9 year LOD signal under the assumption that the latter is caused by the MICG mode. We first note that the inferred amplitude of inner core oscillations, if part of the MICG mode, would then imply similar fluctuations in zonal core flows inside the TC, and these are much larger than the approximate upper bound of $0.02^{\circ}$ stated above. It is possible that zonal core flows are not sufficiently well resolved inside the TC, so their amplitude may be larger. Alternatively, a part of the seismic signal could be explained by changes in the ICB topography (wen06; yao19), including from viscous relaxation (vidale25). In any case, its large amplitude notwithstanding, our results indicate that subdecadal changes in $\Omega_{i}$ cannot be associated with the MICG mode, contrary to the interpretation made by wang22. Instead, we favour an interpretation of subdecadal $\Omega_{i}$ changes entrained by TO modes or, more generally, by Alfvén waves. In this case, unlike for the MICG mode, a prediction of $\Omega_{i}$ cannot be constructed directly from the LOD signal but instead depends on many factors, including the relative strengths of the EM and gravitational torque on the mantle and the structure of Alfvén waves. It is of course possible that the phase of $\Omega_{i}$ entrained by Alfvén waves is similar to that predicted from the LOD signal on the basis of the MICG mode.
Finally, we have noted that the MICG period can act as a resonance to amplify TO modes. More generally any driven motion of the TC at a period that approaches that of the MICG would be amplified. If $\overline{\Gamma}\approx 3\times 10^{19}$ N m, the MICG period is approximately 20 years. If the inner core viscous relaxation time is of the same order or longer than this period, some of the observed multidecadal LOD changes caused by the core (e.g. finlay23) may then result from an amplification of driven zonal core flow motion (e.g. more18) due to their close temporal proximity to the MICG period.
Acknowledgements.
This work was financially supported by a Discovery Grant from NSERC/CRSNG (Canada). Figures were produced using the GMT software (gmt).
Data availability
The codes necessary to reproduce all of our numerical experiments are freely accessible at the following data repository: dumberry25data.
Competing interests
The author has no competing interests.
[…]

### v2

@@
 The mantle-inner core gravitational mode of oscillation in a strong magnetic field regime
-Abstract The mantle-inner core gravitational (MICG) mode is the free mode axial oscillation between the mantle and inner core sustained by the gravitational torque between their degree 2 order 2 density structures.
-Here, we investigate how the MICG mode is affected by oscillations of cylindrical surfaces in the fluid outer core in the form of Alfvén waves.
-The latter are triggered by oscillations of the tangent cylinder (TC) moving jointly with the inner core and propagate away from the rotation axis.
-We show that the MICG mode remains a distinct normal mode of oscillation of the core-mantle system only when the triggered Alfvén waves are attenuated before they traverse the width of the fluid core.
+Abstract The mantle-inner core gravitational (MICG) mode is the free mode of axial oscillation between the mantle and inner core, sustained by the gravitational torque between their degree 2 order 2 density structures.
+As part of this mode, the tangent cylinder (TC) is entrained to move jointly with the inner core, and the oscillations of the TC launch Alfvén waves propagating in the region outside the TC.
+Here, we investigate how the MICG is altered by the strength of the internal magnetic field in the core which controls the travelling speed of Alfvén waves.
+We show that the MICG mode remains a distinct normal mode of oscillation of the core-mantle system only when Alfvén waves are attenuated before they traverse the width of the fluid core.
 For an internal magnetic field strength of a few mT, as we expect in Earth’s core, Alfvén waves can readily traverse the width of the core, and the MICG mode is absorbed into the spectrum of torsional oscillation (TO) modes.
-The MICG period retains a dynamical influence, acting as a point of resonance for TO modes, and marking the transition from a TO mode in which the motion of the TC (including the inner core) is weakly impacted by gravitational coupling to one in which the oscillating motion of the TC is strongly restricted.
+The MICG period retains a dynamical influence, acting as a point of resonance for TO modes, and marking the transition from a TO mode that is weakly impacted by gravitational coupling to one in which the motion of the TC (including the inner core) is strongly restricted.
 Our results imply that the observed 6-year periodic signal in the length of day cannot be interpreted as the signature of the MICG mode and must instead be caused by TO modes, or more generally, by the propagation of Alfvén waves.
-{summary}Non-technical summary The Earth’s mantle is not uniform but contains continental size density anomalies.
+{summary}
+Non-technical summary The Earth’s mantle is not uniform but contains continental size density anomalies.
 The topography of the solid inner core (located at the centre of our planet) is aligned on average with these mantle density anomalies.
@@
 Here, we show that when the magnetic field inside the fluid core is high (of the order of a few mT), the nature of the MICG changes; it is no longer a distinct, independent mode of oscillation of the core-mantle system.
-Instead, the MICG mode merges into the set of modes of torsional oscillations (TO), which consist in the longitudinal oscillations of nested cylindrical surfaces within the fluid core maintained by the magnetic tension between them (oscillations that also known as Alfvén waves).
+Instead, the MICG mode merges into the set of modes of torsional oscillations (TO), which consist in the longitudinal oscillations of nested cylindrical surfaces within the fluid core maintained by the magnetic tension between them (oscillations that are also known as Alfvén waves).
 The inner core can still oscillate with respect to the mantle, but it is due to its entrainment as part of the TO modes.
-The MICG mode retains a dynamical influence, as its period determines whether the inner core oscillation for a given TO mode is impeded or not by the mantle gravitational torque.
 These results have important implications for the interpretation of the 6 year LOD signal, which cannot be caused by the MICG mode, but must instead be caused by the TO modes, or more generally, by the propagation of Alfvén waves in the fluid core.
 1 Introduction
-The mantle-inner core gravitational (MICG) mode of axial oscillation owes its existence to aspherical mass anomalies within the mantle, including those caused by the topography at surfaces of density contrast (buffett96a).
-The largest of these occur at spherical harmonic degree 2 order 2 (e.g.
-defraigne96; simmons07), warping equipotential surfaces everywhere inside the Earth (including the surface geoid) into ellipsoid shapes elongated in the equatorial plane.
+The mantle-inner core gravitational (MICG) mode of axial oscillation owes its existence to aspherical mass anomalies within the mantle, including those caused by the topography at surfaces of density contrast [7].
+The largest of these occur at spherical harmonic degree 2 order 2 [16, 46, e.g.], warping equipotential surfaces everywhere inside the Earth (including the surface geoid) into ellipsoid shapes elongated in the equatorial plane.
 Within the core, surfaces of constant density coincide with these elliptically shaped equipotential surfaces in order to maintain hydrostatic equilibrium.
@@
 Its shape should then be elliptical in the plane of the equator (i.e.
-equatorially elliptical) and longitudinally aligned with the leading order mantle mass anomalies (buffett96a).
-An axial rotation of the inner core out of this longitudinal alignment, driven for example by magnetic stresses at the ICB (e.g.
-buffett00; aubert11), leads to a gravitational torque acting to restore the equilibrium and state of minimum gravitational potential energy (buffett96a).
-This gravitational torque can thus sustain a natural mode of axial oscillation between the inner core and the mantle, referred to as the MICG mode (mound06).
-The period of the MICG mode depends on the gravitational strength factor, $\overline{\Gamma}$, capturing the amplitude of the gravitational torque for a given small longitudinal misalignment between the inner core and mantle (buffett96b).
-In turn, $\overline{\Gamma}$ depends on the amplitude of the degree 2 order 2 geoid at the core-mantle boundary (CMB), which itself depends on estimates of the mantle density field, including that from the topography of the CMB (buffett96a; davies14; chao17; shih21).
-The period of the MICG mode also depends on whether the region of the fluid core inside the axial cylinder that encloses the inner core – the so-called tangent cylinder (TC) – is entrained into co-rotation with the inner core (see eqs.
-(8) and (9) below).
+equatorially elliptical) and longitudinally aligned with the leading order mantle mass anomalies [7].
+An axial rotation of the inner core out of this longitudinal alignment, driven for example by magnetic stresses at the ICB [9, 2, e.g.], leads to a gravitational torque acting to restore the equilibrium state of minimum gravitational potential energy [7].
+This gravitational torque can thus sustain a natural mode of axial oscillation between the inner core and the mantle, referred to as the MICG mode [39].
+The period of the MICG mode depends on the gravitational strength factor, $\overline{\Gamma}$, capturing the amplitude of the gravitational torque for a given small longitudinal misalignment between the inner core and mantle [6].
+In turn, $\overline{\Gamma}$ depends on the amplitude of the degree 2 order 2 geoid at the core-mantle boundary (CMB), which itself depends on estimates of the mantle density anomalies, including that caused by the topography of the CMB [7, 15, 12, 45].
+The period of the MICG mode also depends on whether the region of the fluid core inside the axial cylinder that encloses the inner core – the so-called tangent cylinder (TC) – is entrained into co-rotation with the inner core, which is expected given the strong electromagnetic coupling at the ICB [6, 37].
 Estimates of the period of the MICG mode can be derived by computing predictions of $\overline{\Gamma}$ on the basis of viscous mantle flow models.
-The latter are built from estimates of the density anomalies within the mantle inferred from seismic tomography and the subsequent viscous deformation of the CMB, the surface and other surfaces of density discontinuities that they cause (e.g.
-hager85; forte94; defraigne96).
-Based on the mantle flow models available at the time, buffett96a estimated a MICG period of approximately 2 to 3 years.
-The more recent survey of davies14 suggests a longer period in the range 7 to 18 years.
-The range of uncertainty reflects our limited knowledge of both the amplitude of the internal mantle density anomalies (primarily those of degree 2, order 2 associated with Large Low Seismic Velocity provinces in the lower mantle, e.g.
-garnero16; mcnamara19) and the viscosity of the lowermost mantle.
+The latter are built from density anomalies within the mantle inferred from seismic tomography and the subsequent viscous deformation of surfaces of density discontinuities that they cause [28, 22, 16, e.g.].
+Based on the mantle flow models available at the time, [7] estimated a MICG period of approximately 2 to 3 years.
+The more recent survey of [15] suggests a longer period in the range 7 to 18 years.
+The range of uncertainty reflects our limited knowledge of both the amplitude of the internal mantle density anomalies [23, 36, primarily those of degree 2, order 2 associated with Large Low Seismic Velocity provinces in the lower mantle, e.g.] and the viscosity of the lowermost mantle.
 An alternate approach is to seek the signature of the MICG mode in the observed changes in the mantle rotation rate measured as changes in the length of day (LOD).
-Indeed, an observed quasi-periodic LOD signal with an amplitude of $\sim 0.13$ ms and a period of $\sim 5.9$ days (abarcadelrio00; holme13; chao14) has been linked with the MICG mode (mound06; chao17; ding18; duan18; shih21).
-The magnitude of $\overline{\Gamma}$ required to match this period provides then a direct measure of the equatorial ellipticities of the geoid at the CMB and of the ICB topography (davies14; chao17; shih21).
+Indeed, an observed quasi-periodic LOD signal with an amplitude of $\sim 0.13$ ms and a period of $\sim 5.9$ years [1, 30, 13] has been linked with the MICG mode [39, 12, 17, 18, 45].
+The magnitude of $\overline{\Gamma}$ required to match this period provides then a direct measure of the equatorial ellipticities of the geoid at the CMB and of the ICB topography [15, 12, 45].
 This simplified picture of the MICG mode neglects the dynamical influence of the fluid core in the region outside the tangent cylinder (ROTC) or assumes that it plays a passive role.
-At periods relevant to the MICG, axisymmetric azimuthal (zonal) flows are expected to be invariant in the direction of the rotation axis and to take the form of rigidly rotating nested cylindrical surfaces (jault08).
+At periods relevant to the MICG mode, axisymmetric azimuthal (zonal) flows are expected to be invariant in the direction of the rotation axis and take the form of rigidly rotating nested cylindrical surfaces [32].
 A differential motion between adjacent cylinders shears the cylindrically radial ($B_{s}$) magnetic field lines that threads their surfaces.
-This creates an azimuthal restoring force between cylinders and allows for the propagation of Alfvén waves propagating in the cylindrically radial direction (braginsky70).
-The normal modes of axial oscillations between the cylinders formed by the superposition of these Alfvén waves are referred to as torsional oscillations (TO, braginsky70) and their periods depend the strength of $B_{s}$ within the core.
-The prevalent view at the time the earliest studies on the MICG mode were published (e.g.
-buffett96a; buffett96b; mound06) was that the gravest mode of TO had a period of 60-80 years, which then implied a typical $B_{s}$-field strength of 0.2-0.3 mT within the core (zatman97).
-Under this scenario, a subdecadal MICG mode has a distinctly shorter period, so the ROTC may be considered to play a passive role even if it is engaged in the MICG to some degree by the propagation of Alfvén waves (e.g.
-mound03; mound06).
+This creates an azimuthal restoring force between cylinders and allows for the propagation of Alfvén waves in the cylindrically radial direction [4].
+The normal modes of axial oscillations between the cylinders formed by the superposition of these Alfvén waves are referred to as torsional oscillations [4, TO,] and their periods depend the strength of $B_{s}$ within the core.
+These TO modes include the motion of the TC, so they are necessarily coupled to the MICG mode.
+The prevalent view at the time the earliest studies on the MICG mode were published [7, 6, 37, 39, e.g.] was that the gravest mode of TO had a period of 60-80 years, implying a $B_{s}$-field strength of 0.2-0.3 mT within the core [51].
+Under this scenario, the MICG period is distinctly shorter than that of the fundamental TO mode.
+Consequently, the MICG mode provides a resonant band of amplification to higher TO harmonics that have periods in the vicinity of the MICG period [37].
+The oscillating motion of the inner core, entrained by the TO modes, exerts a gravitational torque on the mantle leading to distinct peaks in the LOD at periods close to the MICG mode.
+Even though the ROTC is dynamically involved, it carries little angular momentum and the primary angular momentum balance for these amplified TO modes is between the mantle and the whole of the TC, just as is the case for a pure MICG mode.
+The 5.9 yr oscillation in the LOD may still be tied to MICG period and provides a constraint on $\overline{\Gamma}$ [39].
 The general consensus on the magnetic field strength within the core has since shifted.
-Numerical models of the dynamo suggest that the field deep inside the core is likely stronger by a factor 10 than that at the CMB, suggesting an internal field strength of 3-4 mT or higher (e.g.
-christensen06; aubert09).
-At such a strength, Alfvén waves should propagate through the entire width of the core in only a few years and the periods of the gravest TO modes are expected to be in the subdecadal range.
-Indeed, gillet10 showed that the angular momentum carried by propagating Alfvén waves in the ROTC matches that required to explain the observed periodic $5.9$-yr LOD signal.
-This latter result indicates that, for the 5.9 yr LOD signal, the mantle exchanges its angular momentum primarily with the ROTC.
-This is in contrast to the oscillation behaviour expected for the MICG mode, where the mantle exchanges its angular momentum primarily with the inner core (whether entraining the entire TC or not) and not the ROTC.
-It also presents an alternative explanation for the 5.9 yr LOD signal, which would then represents the entrainment of the mantle into oscillation as a result of these propagating Alfvén waves.
-Or, at the very least, if the MICG mode remains connected to the 5.9 yr LOD signal, the ROTC is not passive and is instead entrained into motion by the propagation of Alfvén waves.
-Here, our goal is to investigate how the MICG mode is altered when a strong $B_{s}$-field inside the core allows for the propagation of fast Alfvén waves that can travel through the width of the core in a timescale shorter or similar to the MICG period.
-We shall show that the MICG mode is no longer an independent oscillation mode of the coupled core-mantle system but rather gets absorbed into the spectrum of TO modes, although the latter may be modified by gravitational coupling.
+Numerical models of the dynamo suggest that the field deep inside the core is likely stronger by a factor 10 than at the CMB, suggesting an internal field strength of 3-4 mT [14, 3, e.g.].
+With such a large field strength, Alfvén waves propagate through the entire width of the core in only a few years and the periods of the gravest TO modes are expected to be in the subdecadal range.
+Indeed, [25] showed that the angular momentum carried by propagating Alfvén waves matches that required to explain the observed periodic $5.9$-yr LOD signal.
+This indicates that the mantle exchanges its angular momentum with the whole of the core rather than with the TC alone, in contrast with an idealized MICG mode.
+Furthermore, in this scenario, the 5.9 yr period is instead tied to the period of the fundamental mode of TO, not the strength of the gravitational torque, even though the gravitational torque may still provide the means by which angular momentum is transferred to the mantle.
+In a strong $B_{s}$-field regime, Alfvén waves travel through the width of the core in a shorter time than the MICG period.
+Here, our goal is to investigate how the MICG mode is altered by such fast propagating Alfvén waves.
+We shall show that when Alfvén waves can propagate through the ROTC before they get attenuated, the MICG mode is no longer an independent mode of the coupled core-mantle system.
+Instead, it gets absorbed into the spectrum of TO modes, although the latter are affected by gravitational coupling.
 2 Theory
 2.1 Angular momentum balance
-The model of internal coupling between the mantle, fluid core and inner core that we use follows broadly that presented in several studies, including in buffett96b; buffett98; mound03; mound05b; dumberry08c; dumberry10b.
+The model of internal coupling between the mantle, fluid core and inner core that we use follows broadly that presented in several studies, including in [6, 8, 37, 38, 20, 21].
 We assume a spherical model of Earth’s interior comprised of an inner core (radius $r_{i}$), a fluid core (outer radius $r_{f}$) and mantle.
@@
 $C_{i}\frac{d}{dt}\Omega_{i}$ $=-\overline{\Gamma}\alpha+\Gamma_{icb}\,$, (1b)
-[…]
+$\int_{0}^{r_{f}}c_{f}\,\frac{d\Omega_{f}}{dt}\,ds$ $=-\Gamma_{cmb}-\Gamma_{icb}\,$, (1c)
 where, $C_{m}$, $C_{i}$ are the axial moments of inertia of the mantle and inner core, $\Omega_{m}$, $\Omega_{i}$ their rotation rates, and
@@
 $\Gamma_{cmb}$ and $\Gamma_{icb}$ are the torques from all surface forces acting on the mantle at the CMB and on the inner core at the ICB, respectively.
-$\overline{\Gamma}$ is the strength of gravitational torque between the mantle and inner core and $\alpha$ is the (assumed small) longitudinal angle of misalignment between the degree 2 order 2 mantle density field and the long equatorial axis of the ICB topography.
+$\overline{\Gamma}$ is the strength of the gravitational torque between the mantle and inner core and $\alpha$ is the (assumed small) longitudinal angle of misalignment between the degree 2 order 2 mantle density field and the long equatorial axis of the ICB topography.
 $\alpha$ depends on the differential rotation between the inner core and the mantle and on the viscous relaxation time $\tau$} for the ICB topography to return to its equilibrium alignment with the imposed gravitational potential from the mantle; its evolution is given by
 $\frac{d\alpha}{dt}=\Omega_{i}-\Omega_{m}-\frac{\alpha}{\tau}\,.$ (3)
-We assume that $\Gamma_{icb}$ and $\Gamma_{cmb}$ are both dominated by electromagnetic (EM) forces, and expressed as (e.g. buffett98)
-[…]
-[…]
+We assume that $\Gamma_{icb}$ and $\Gamma_{cmb}$ are both dominated by electromagnetic (EM) forces, and expressed as [8, e.g.]
+$\Gamma_{cmb}$ $=\int_{0}^{r_{f}}{\cal F}_{m}(s)\left[\Omega_{f}-\Omega_{m}\right]ds\,$, (4a)
+$\Gamma_{icb}$ $=\int_{0}^{r_{i}}{\cal F}_{i}(s)\left[\Omega_{f}-\Omega_{i}\right]ds\,$, (4b)
 where the coupling parameters ${\cal F}_{m}(s)$ and ${\cal F}_{i}(s)$ depend on the morphology of the magnetic field and the electrical conductivity of the solid side of the boundary.
@@
 ${\cal F}_{i}(s)$ $=4\pi s^{3}G_{i}\frac{z_{i}}{r_{i}}\left(B_{i}^{d}\right)^{2}\,.$ (5b)
-where $B_{m}^{d}$ and $B_{i}^{d}$ are the polar amplitudes of the axial dipole at the CMB and ICB, respectively, $\langle B_{r,m}\rangle$ is the root mean squared (rms) strengths of the total radial field (including the axial dipole) at the CMB, $G_{m}$ is the conductance of the lowermost mantle and $G_{i}$ the conductance factor at the ICB given by
+where $B_{m}^{d}$ and $B_{i}^{d}$ are the polar amplitudes of the axial dipole at the CMB and ICB, respectively, $\langle B_{r,m}\rangle$ is the root mean squared (rms) strength of the total radial field (including the axial dipole) at the CMB, $G_{m}$ is the conductance of the lowermost mantle and $G_{i}$ is a conductance factor at the ICB given by
 $G_{i}=\frac{1}{4}[1+i\,\mbox{sgn}(\omega)]\,\sigma\,\delta\,$, (6)
 where $\sigma$ is the conductivity of the fluid core (assumed equal to that of the inner core), $\delta=\sqrt{2/|\omega|\mu\sigma}$ is the magnetic skin depth at oscillation frequency $\omega$, with $\mu$ the magnetic permeability of free space, and $\mbox{sgn}(\omega)=\omega/|\omega|$.
-The form of ${\cal F}_{m}(s)$ was shown in dumberry08c to be a good match with the EM coupling resulting from a realistic field model at the CMB comprised of a large number of spherical harmonic terms.
+The form of ${\cal F}_{m}(s)$ was shown in [20] to be a good match with the EM coupling resulting from a realistic field model at the CMB comprised of a large number of spherical harmonic terms.
 The field at the ICB is undoubtedly more complex than a simple axial dipole, although the latter is expected to be its dominant contribution, and it is sufficient to capture the first order effect of EM coupling at the ICB.
@@
 To illustrate how the MICG mode emerges from the above system, let us assume a perfectly rigid inner core ($\tau=\infty$), and no coupling at the CMB (${\cal F}_{m}(s)=0$).
-Let us further assume that the fluid inside the TC rotates as a single rigid body with angular velocity $\Omega_{f}^{tc}$.
+Let us further assume that the fluid in the ROTC is not involved and remains at rest, while the fluid inside the TC rotates as a single rigid body with angular velocity $\Omega_{f}^{tc}$.
 Taking the time-derivatives of eqs.
@@
 (3), gives
-$C_{m}\frac{d^{2}\Omega_{m}}{dt^{2}}$ $=\bar{\Gamma}(\Omega_{i}-\Omega_{m})\,,$ (7a)
-$C_{i}\frac{d^{2}\Omega_{i}}{dt^{2}}$ $=-\bar{\Gamma}(\Omega_{i}-\Omega_{m})+r_{i}\,\overline{\cal F}_{i}\frac{d}{dt}(\Omega_{f}^{tc}-\Omega_{i})\,,$ (7b)
+$C_{m}\frac{d^{2}\Omega_{m}}{dt^{2}}$ $=\overline{\Gamma}(\Omega_{i}-\Omega_{m})\,$, (7a)
+$C_{i}\frac{d^{2}\Omega_{i}}{dt^{2}}$ $=-\overline{\Gamma}(\Omega_{i}-\Omega_{m})+r_{i}\,\overline{\cal F}_{i}\frac{d}{dt}(\Omega_{f}^{tc}-\Omega_{i})\,$, (7b)
 $C_{f}^{tc}\frac{d^{2}\Omega_{f}^{tc}}{dt^{2}}$ $=-r_{i}\,\overline{\cal F}_{i}\frac{d}{dt}(\Omega_{f}^{tc}-\Omega_{i})\,$, (7c)
-where $C_{f}^{tc}$ is the moment of inertia of the fluid inside the TC and ¯$\overline{\cal F}_{i}$ is the integrated average of ${\cal F}_{i}(s)$ over the ICB.
+where $C_{f}^{tc}$ is the moment of inertia of the fluid inside the TC and $\overline{\cal F}_{i}$ is the integrated average of ${\cal F}_{i}(s)$ over the ICB.
 This forms a simple system of axial oscillations between $\Omega_{m}$, $\Omega_{i}$ and $\Omega_{f}^{tc}$, with a restoring gravitational torque between the inner core and mantle and a restoring (and dissipating) EM torque between the inner core and the fluid inside the TC.
-Note that the fluid in the ROTC is not involved and assumed to remain at rest.
-The natural modes of oscillations of this system include the MICG mode, whose frequency depends on the strength of $\overline{\Gamma}$ and ¯$\overline{\cal F}_{i}$.
-Let us consider two end-member scenarios for ¯$\overline{\cal F}_{i}$.
-For no EM coupling at the ICB (¯$\overline{\cal F}_{i}=0$), the fluid inside the TC is decoupled and not involved in the MICG mode.
-The frequency and eigenvector of the MICG mode are (dumberry10b),
-$\omega_{micg}=\sqrt{\frac{\bar{\Gamma}(C_{m}+C_{i})}{C_{m}C_{i}}}\,,\hskip 28.45274pt\left[\Omega_{m},\Omega_{i}\right]=\left[1,-\frac{C_{m}}{C_{i}}\right]\,.$ (8)
-The oscillation (and angular momentum exchange) is between the mantle and inner core, with no involvement of any parts of the fluid core.
-As an example, taking $\overline{\Gamma}=3\times 10^{20}$ N m, and $C_{m}$ and $C_{i}$ given Table 1, this gives a MICG period of 2.8 years.
-If instead ¯$\overline{\cal F}_{i}$ is very large, the differential rotation between $\Omega_{f}^{tc}$ and $\Omega_{i}$ must remain very small (over the period of the MICG mode), and $\Omega_{f}^{tc}$ and $\Omega_{i}$ are locked into a common rotation.
+The natural modes of oscillations of this system include the MICG mode, whose frequency depends on the strength of $\overline{\Gamma}$ and $\overline{\cal F}_{i}$.
+Let us consider two end-member scenarios for $\overline{\cal F}_{i}$.
+For no EM coupling at the ICB ($\overline{\cal F}_{i}=0$), the fluid inside the TC is decoupled and not involved in the MICG mode.
+The frequency and eigenvector of the MICG mode are [21, e.g.],
+$\omega_{micg}=\sqrt{\frac{\bar{\Gamma}(C_{m}+C_{i})}{C_{m}C_{i}}}\,,\hskip 17.07182pt\left[\Omega_{m},\Omega_{i}\right]=\left[1,-\frac{C_{m}}{C_{i}}\right]\,.$ (8)
+The oscillation (and angular momentum exchange) is between the mantle and inner core, with no involvement of the fluid core.
+As an example, taking $\overline{\Gamma}=3\times 10^{20}$ N m, and $C_{m}$ and $C_{i}$ given in Table 1, this gives a MICG period of 2.8 years.
+If instead $\overline{\cal F}_{i}$ is very large, the differential rotation between $\Omega_{f}^{tc}$ and $\Omega_{i}$ must remain very small (over the period of the MICG mode), and $\Omega_{f}^{tc}$ and $\Omega_{i}$ are locked into a common rotation.
 Substituting the EM coupling term of eq.
@@
 (7b), and assuming $\Omega_{f}^{tc}\approx\Omega_{i}$, the frequency and eigenvector of the MICG mode are now
-[…]
+$\omega_{micg}=\sqrt{\frac{\bar{\Gamma}(C_{m}+C_{tc})}{C_{m}C_{tc}}}\,,\hskip 17.07182pt\left[\Omega_{m},\Omega_{i}\right]=\left[1,-\frac{C_{m}}{C_{tc}}\right]\,$, (9)
 where $C_{tc}=C_{i}+C_{f}^{tc}$.
@@
 Using $\overline{\Gamma}=3\times 10^{20}$ N m as above, the period is now 5.9 years, approximately twice as long.
-If ¯$\overline{\cal F}_{i}$ is somewhere between these two extremes, with an amplitude such that the time it takes for $\Omega_{i}$ and $\Omega_{f}^{tc}$ to be brought back into co-rotation is of the order a few years, then the MICG mode is more complex, with each of $\Omega_{m}$, $\Omega_{i}$ and $\Omega_{f}^{tc}$ oscillating at different phases.
-At periods of a few years, and assuming a radial magnetic field strength at the ICB of the order of 1 mT or larger, EM coupling is sufficiently strong to prevent any large differential rotation at the ICB (gubbins81).
+If $\overline{\cal F}_{i}$ is somewhere between these two extremes, with an amplitude such that the time it takes for $\Omega_{i}$ and $\Omega_{f}^{tc}$ to be brought back into co-rotation is of the order a few years, then the MICG mode is more complex, with each of $\Omega_{m}$, $\Omega_{i}$ and $\Omega_{f}^{tc}$ oscillating at different phases.
+At periods of a few years, and assuming a radial magnetic field strength at the ICB of the order of 1 mT or larger, EM coupling is sufficiently strong to prevent any large differential rotation at the ICB [27].
 A rotation of the inner core will tend to entrain the fluid in contact with the ICB into co-rotation.
-This change in zonal flow near the ICB is communicated by the (strong) Coriolis force to the whole of the fluid column above and below the inner core, forcing it to rotate rigidly with it in accordance with the Proudman-Taylor constraint (jault08).
+This zonal motion is communicated by the (strong) Coriolis force to the whole of the fluid column above and below the inner core, forcing it to rotate rigidly with it in accordance with the Proudman-Taylor constraint [32].
 Hence, on dynamical grounds, and for the expected strength of the magnetic field at the ICB, we expect that the MICG mode involves the whole of the TC, not just the inner core.
 This simple presentation of the MICG mode is idealized in several ways.
-First, although ¯$\overline{\cal F}_{i}$ is most likely strong enough to lock the fluid cylinders inside the TC to the rotational motion of the inner core, some degree of differential rotation can occur near the outer edge of the TC (e.g.
-buffett96b; buffett98).
+First, although $\overline{\cal F}_{i}$ is most likely strong enough to lock the fluid cylinders inside the TC to the rotational motion of the inner core, some degree of differential rotation can occur near the outer edge of the TC [6, 8, e.g.].
 Second, we have neglected the global elastic deformations that occur in response to the rotational displacement of the inner core out of its equilibrium longitudinal alignment.
 Taking these into account would lengthen the MICG period, though not by a large fraction.
-Third, if the inner core relaxes viscously over a relatively short time (approaching the MICG period), not only this can also lengthen the MICG period, it can lead to a significant attenuation of the mode.
-A fourth way this discussion of the MICG mode is idealized is that it assumes a complete decoupling of the fluid in the ROTC.
-Indeed, assessing the role of the ROTC on the MICG mode is the motivation for our study.
+Third, if the inner core relaxes viscously over a relatively short time (approaching the MICG period), this would attenuate the mode significantly.
+Perhaps the most important dynamical omission from this idealized picture though is the role played by the fluid in the ROTC.
+The latter has been considered in a number of studies [37, 38, 10, 39, e.g.], although under the assumption that the magnetic field permeating the entire core is relatively weak, with a strength of the order of 0.3 to 0.4 mT.
+The primary motivation for our study is precisely to reassess how the dynamics of the ROTC affects the MICG mode when the magnetic field inside the core is much stronger.
 2.3 Torsional oscillation modes
 The dynamical equation that governs the evolution of $\Omega_{f}$ is obtained by integrating the azimuthal component of the momentum equation over a cylindrical surface in the fluid core.
-For this specific geometry, the Coriolis acceleration term vanishes (e.g.
-taylor63; braginsky70), giving
-$c_{f}\frac{\partial\Omega_{f}}{\partial t}=\frac{1}{\rho\mu}\frac{\partial}{\partial s}\left(c_{f}\{B_{s}\}\frac{b_{\phi}}{s}\right)-{\cal F}_{m}\left[\Omega_{f}-\Omega_{m}\right]-{\cal F}_{i}\left[\Omega_{f}-\Omega_{i}\right]+\nu\frac{\partial}{\partial s}\left(c_{f}\frac{\partial\Omega_{f}}{\partial s}\right)\,,$ (10)
+For this specific geometry, the Coriolis acceleration term vanishes [47, 4, e.g.], giving
+[…]
 where $\nu$ is the kinematic viscosity, $\{B_{s}\}$ is the rms strength of the $s$-component of the background magnetic field averaged over a cylindrical surface, and $b_{\phi}$ is the azimuthal magnetic field perturbation induced by the differential motion of the cylinders.
@@
 (11) into eq.
-(10), we retrieve the characteristic equation for Alfvén waves subject to EM friction at the top and bottom of the cylinders, (e.g.
-braginsky70; buffett98)
-$c_{f}\frac{\partial^{2}\Omega_{f}}{\partial t^{2}}=\frac{1}{\rho\mu}\frac{\partial}{\partial s}\left(c_{f}\{B_{s}^{2}\}\frac{\partial\Omega_{f}}{\partial s}\right)-{\cal F}_{m}\frac{\partial}{\partial t}\left[\Omega_{f}-\Omega_{m}\right]-{\cal F}_{i}\frac{\partial}{\partial t}\left[\Omega_{f}-\Omega_{i}\right]\,,$ (12)
+(10), we retrieve the characteristic equation for Alfvén waves subject to EM friction at the top and bottom of the cylinders, [4, 8, e.g.]
+[…]
 where we have assumed $\{B_{s}\}^{2}\approx\{B_{s}^{2}\}$.
 The normal modes of oscillations of $\Omega_{f}$ that obey eq.
-(12) are the torsional oscillations (TO) modes (e.g.
-braginsky70).
+(12) are the torsional oscillations (TO) modes [4, e.g.].
 In their simplest form, when assuming ${\cal F}_{m}={\cal F}_{i}=0$, these modes describe the set of azimuthal oscillations of cylinders maintained by the magnetic tension between them.
 The terms that contain ${\cal F}_{m}$ and ${\cal F}_{i}$ capture the coupling of the fluid cylinders with the mantle (at the CMB) and the inner core (at the ICB).
-Keeping these terms not only provides a way by which TO modes can exchange angular momentum with the mantle and inner core, it also influences the structure of the modes.
-In all previous studies on TO modes coupled with the mantle and inner core, eq.
-(12) was assumed for the dynamics of $\Omega_{f}$ (buffett96b; buffett98; mound03; mound05b; dumberry08c; dumberry10b).
+Keeping these terms affects the structure of the modes and provides a way by which TO modes can exchange angular momentum with the mantle and inner core.
+In most previous studies on TO modes coupled with the mantle and inner core, eq.
+(12) was assumed for the dynamics of $\Omega_{f}$ [6, 8, 37, 38, 20, 21].
+The studies of [5] and [40] included the effects of viscous diffusion both within the volume of the fluid core and at boundaries.
 Here, we retain both magnetic and viscous diffusion and solve for both $\Omega_{f}$ and $b_{\phi}$ using the coupled eqs.
 (10) and (11).
-The motivation for retaining magnetic diffusion is grounded in the results of luo22a who showed that it reduces the quality factor of TO modes to values below 10.
 Although we expect $\nu\ll\eta$ in the fluid core, the motivation for retaining viscous diffusion in (10) is for numerical convenience, as it makes it easier to prescribe appropriate boundary conditions for $\Omega_{f}$.
 For all our results, we have used $\nu=\eta/100$, so as to ensure that magnetic diffusion dominates over viscous diffusion.
+Although it would be preferable to solve for $\Omega_{f}$ and $b_{\phi}$ as a function of both $s$ and $z$ [35, e.g.], our simple one-dimensional (1D) model captures the leading order dynamics of Alfvén waves within the core.
+As we shall show, whether the MICG mode is an independent mode of the core-mantle system depends on the travel time versus diffusion time of Alfvén waves in the ROTC.
+Alfvén waves travel at velocity $V_{A}=\{B_{s}\}/\sqrt{\rho\mu}$, and the time they take to travel between the TC and the equator of the CMB ($s=r_{f}$) is $\tau_{A}=(r_{f}-r_{i})/V_{A}$.
+Their typical timescale of diffusion from ohmic dissipation within the core is $\tau_{d}=1/(k_{s}^{2}\eta)$, where $k_{s}$ is the wavenumber in the $s$-direction.
+Alfvén waves obey the dispersion relation $\omega=k_{s}V_{A}$, so TO modes that have frequencies close to MICG mode should have a typical wavenumber $k_{s}=\omega_{micg}/V_{A}$, giving $\tau_{d}=V_{A}^{2}/\eta\omega_{micg}^{2}$.
+Alfvén waves are also attenuated by EM coupling at the CMB.
+A simple estimate of this EM attenuation timescale can be obtained from eq.
+(10) and is $\tau_{em}=c_{f}/{\cal F}_{m}$.
+This quantity depends on $s$ but a rough bulk estimate is given by $\tau_{em}=\rho r_{c}/(G_{m}\langle B_{r,m}\rangle^{2})$.
+The travel time versus diffusion time of Alfvén waves is then characterized by two Lundquist numbers,
+$Lu=\frac{\tau_{d}}{\tau_{A}}\,,\qquad Lu_{em}=\frac{\tau_{em}}{\tau_{A}}\,.$ (13)
+If either $Lu<1$ or $Lu_{em}<1$, Alfvén waves are attenuated before they travel the full width of the ROTC.
+Another useful diagnostic quantity is the Lundquist number based on the travel versus diffusion time of Alfvén waves in the conducting mantle,
+$Lu_{m}=\sqrt{\frac{\mu}{\rho}}\,G_{m}\langle B_{r,m}\rangle\,.$ (14)
+In particular, $Lu_{m}$ determines the degree of reflection of Alfvén waves when they reach the equator of the CMB; when $Lu_{m}\approx 1$, Alfvén waves are fully absorbed [44, 24].
+This quantity is denoted as $Q$ in both [44] and [24], but here we reserve $Q$ to denote the quality factor of modes.
 3 Numerical implementation
@@
 For all our results, we used $\zeta=0.001$.
-Equations (10) and (11) each require two boundary conditions, one each at $s_{o}$ and $s_{e}$.
+Eqs.
+(10) and (11) each require boundary conditions at $s_{o}$ and $s_{e}$.
 At $s=s_{o}$, we impose the regularity conditions $d\Omega_{f}/ds=0$ and $b_{\phi}=0$.
-At $s=s_{e}$, we set $d\Omega_{f}/ds=0$ which implies that there is no viscous torque between the fluid core and mantle and minimizes the influence of viscosity on our solutions.
-The condition on $b_{\phi}$ at $s_{e}$ is set by the magnetic field perturbation resulting from EM coupling (buffett98)
-$b_{\phi}(s_{e})=-B_{s_{e}}\mu G_{m}s_{e}\left(\Omega_{f}(s_{e})-\Omega_{m}\right)\,,$ (14)
-where $B_{s_{e}}$ is the rms amplitude of the radial magnetic field at the equator of the CMB.
+At $s=s_{e}$, the condition on $b_{\phi}$ is set by the magnetic field perturbation resulting from EM coupling [8, e.g.
+eq.
+12 of]
+$b_{\phi}(s_{e})=-\langle B_{r,m}\rangle\mu G_{m}s_{e}\left(\Omega_{f}(s_{e})-\Omega_{m}\right)\,$, (16)
+and we impose a free-slip condition, $d\Omega_{f}/ds=0$.
+The latter implies that there is no viscous torque between the fluid core and mantle and minimizes the influence of viscosity on our solutions.
+Since we use $\nu=\eta/100$, the free-slip condition is met by a flow adjustment in a thin viscous boundary layer near $s=s_{e}$.
 With these boundary conditions, integrating eq. (10) between $s_{o}$ and $s_{e}$ yields
-[…]
+$\int_{s_{o}}^{s_{e}}c_{f}\frac{d\Omega_{f}}{dt}ds=-\hat{\Gamma}_{cmb}-\hat{\Gamma}_{icb}\,$, (17)
 where $\hat{\Gamma}_{cmb}$ and $\hat{\Gamma}_{icb}$ are modified expressions for the torque on the mantle and inner core given by
-[…]
-[…]
+$\hat{\Gamma}_{cmb}$ $=\int_{s_{o}}^{s_{e}}{\cal F}_{m}(s)\left[\Omega_{f}-\Omega_{m}\right]ds$
+$\hskip 14.22636pt+4\pi s_{e}^{3}\sqrt{1-s_{e}^{2}}\,G_{m}\langle B_{r,m}\rangle^{2}\left(\Omega_{f}(s_{e})-\Omega_{m}\right)\,$, (18a)
+$\hat{\Gamma}_{icb}$ $=\int_{s_{o}}^{s_{tc}}{\cal F}_{i}(s)\left[\Omega_{f}-\Omega_{i}\right]ds\,.$ (18b)
 The same modified expressions are used for the equations of the mantle (1a) and inner core (1b) to ensure that angular momentum is conserved.
@@
 The crowding of points near $s_{o}$, $s_{tc}$ and $s_{e}$ helps to resolve the thin viscous and magnetic boundary layers at $s_{o}$ and $s_{e}$ and to handle the rapidly changing $c_{f}$ in the region near the TC.
-We use 350 points inside the TC and 650 in the ROTC, for a total of $N_{s}$=1000 grid points, which proved to be sufficient to resolve the structure of the TO modes, and to ensure that angular momentum is conserved to within approximately 1 part in $10^{4}$.
+We use 350 points inside the TC and 650 in the ROTC, for a total of $N_{s}$=1000 grid points, which proved to be sufficient to resolve the structure of the relevant TO modes, and to ensure that angular momentum is conserved to within approximately 1 part in $10^{4}$.
 The set of eqs.
@@
 where $\mathsf{A}$} and $\mathsf{B}$} are $n\times n$ matrices.
-The resulting set of modes separate into a set of $i\omega$ with negative real part (decay) and a complex conjugate imaginary part (frequency); we only solve for the positive frequencies $\omega$.
+The solutions are the possible free modes of the system.
+The imaginary part of $\omega$ is negative, capturing mode decay.
+The real part of $\omega$ is either positive or negative (with the same magnitude) and represents the frequency of oscillation; we only solve for the positive frequencies.
 Note that ${\cal F}_{i}$ in eq.
 (5b) has a weak dependence on $\omega$ through the magnetic skin depth $\delta$.
-However, for a radial magnetic field strength at the ICB of 1 mT and above, the fluid cylinders inside the TC become quasi-locked to the inner core, and this frequency dependence only has a very small effect on the normal modes.
-To simplify, we use a nominal frequency of $62\pi/$ 6 yr-1 for calculating $\delta$, such that ${\cal F}_{i}$ is no longer a function of $\omega$, allowing us to write the system in the simple form of eq.
+However, for a radial magnetic field strength at the ICB $\geq 1$ mT, the fluid cylinders inside the TC become quasi-locked to the inner core, and this frequency dependence only has a very small effect on the normal modes.
+To simplify, we use a nominal frequency of $2\pi/6$ yr-1 for calculating $\delta$, such that ${\cal F}_{i}$ is no longer a function of $\omega$, allowing us to write the system in the simple form of eq.
 (19).
-The same strategy has been employed in previous studies (e.g.
-buffett98; buffett05).
+The same strategy has been employed in previous studies [8, 10, e.g.].
 Solutions depend on the choice of the profile of $\{B_{s}\}$ inside the core. We use a simple prescription given by
-$\{B_{s}\}=\overline{B}_{s}\sin\left(\frac{\pi(s-s_{o})}{s_{e}-s_{o}}\right)+\left(\frac{s-s_{o}}{s_{e}-s_{o}}\right)\frac{\overline{B}_{s}}{6}\,.$ (18)
-The $\sin$ term produces a maximum field of $\overline{B}_{s}$ in the bulk of the region outside the TC, similar to the $B_{s}$ profile inverted by gillet10.
-The linear term ensures that the radial field at the equator does not vanish, and is set such that $B_{s_{e}}=\overline{B}_{s}/6$.
-The numerical values that we adopt for all other parameters are given Table 1.
-The dipole field amplitude $B_{m}^{d}=0.319$ mT and the total rms strength of the radial field $\langle B_{r,m}\rangle=0.391$ mT at the CMB were obtained by taking the time-average of the gufm1 model of jackson00 over the time period 1950–1990 (the rms of the radial part of the dipole field alone is 0.226 mT).
-Table 1: Parameters used in calculations
+$\{B_{s}\}=\overline{B}_{s}\sin\left(\frac{\pi s}{r_{f}}\right)+\left(\frac{s}{r_{f}}\right)B_{s_{e}}\,$, (20)
+where $\overline{B}_{s}$ is an amplitude factor and $B_{s_{e}}$ is the amplitude of the radial magnetic field at the equator of the CMB.
+The sine term produces a maximum $B_{s}$-field in the middle of the core, similar to the profile inverted by [25].
+The linear term ensures that the radial field at the equator does not vanish.
+For our numerical experiments in the weak-field regime, $\overline{B}_{s}$ is varied between 0.1-0.3 mT, and we set $B_{s_{e}}=\overline{B}_{s}/6$.
+In the strong-field regime, we set $\overline{B}_{s}=3.3$ mT and $B_{s_{e}}=\langle B_{r,m}\rangle$.
+The numerical values that we adopt for all other parameters are given in Table 1.
+The dipole field amplitude $B_{m}^{d}=0.319$ mT and the total rms strength of the radial field $\langle B_{r,m}\rangle=0.391$ mT at the CMB were obtained by taking the time-average of the gufm1 model of [31] over the time period 1950–1990 (the rms of the radial part of the dipole field alone is 0.226 mT).
+We note that, for our experiments in the weak-field regime, there is an inconsistency between our chosen ICB and CMB field strengths and the lower ${B}_{s}$-field in the range of 0.1-0.3 mT.
+These low values were chosen because they are similar to those used in previous studies of the MICG mode [37, 39, e.g.] and they help to illustrate the role that Alfvén waves have on this mode.
+We show results for different choices of mantle conductance: $G_{m}=10^{8}$ S, $2\times 10^{8}$ S, $4\times 10^{8}$ S, and $6\times 10^{8}$ S.
+The corresponding $Lu_{m}$ values, in increasing order of $G_{m}$, are: 0.34, 0.68, 1.36, 2.05.
+These remain sufficiently low that our 1D model captures correctly the reflection properties of Alfvén waves as they reach the equator of the CMB [44, 24].
+Table 1: Parameters used in calculations.
 Parameter value
@@
 axial dipole, ICB $B_{i}^{d}=3.0$ mT
-(*) Based on a uniform fluid core density of $\rho=1.1\times 10^{4}$ kg m-3
+(*) Based on a uniform fluid core density of $\rho=1.1\times 10^{4}$ kg m-3.
 4 Results
 Let us first illustrate a case where the modes of axial oscillations between the mantle, inner core and fluid core include a distinct MICG mode.
-We set $\overline{\Gamma}=3\times 10^{20}$ N m, which gives an expected theoretical MICG frequency of $\omega_{micg}=1.065$ yr-1 (period of 5.902 years) based on eq.
+We set $\overline{\Gamma}=3\times 10^{20}$ N m, which gives an expected MICG frequency of $\omega_{micg}=1.065$ yr-1 (period of 5.902 years) based on eq.
 (9), and we set the mantle conductance at $G_{m}=10^{8}$ S.
 We set $\tau=100$ years to ensure that the MICG mode is only weakly damped by the viscous relaxation of the inner core.
-A key feature of the MICG mode is a weak involvement of the ROTC and this can be achieved by choosing a sufficiently weak $B_{s}$-field amplitude across the TC; we take $\overline{B}_{s}=0.1$ mT.
-We then compute the set of normal modes of oscillations for this setup, which we refer to as case $a$.
-Figure 1a shows the frequency and $Q$ of a subset of the resulting modes (light blue circles) that have frequencies in the vicinity of $\omega_{micg}$.
+We choose $\overline{B}_{s}=0.1$ mT, which gives a typical travel time for Alfvén waves across the ROTC of $\tau_{A}=84$ yr.
+We compute the set of normal modes of oscillations for this setup, which we refer to as case $a$.
+Figure 1a shows the frequency and quality factor $Q$ of a subset of the resulting modes (light blue circles) that have frequencies in the vicinity of $\omega_{micg}$.
 For this choice of $\overline{B}_{s}$, the fundamental mode (first harmonic) of TO has a period of $\sim 180$ years; the sequence of modes that have a $Q$ between 3 and 4 on Figure 1a correspond to higher harmonics of TO.
@@
 Figure 2a shows the profile of $\Omega_{f}$ for this MICG mode, at the phase when $\Omega_{i}$ is maximum in the prograde direction.
-Inside the TC, $\Omega_{f}$ is ocked with $\Omega_{i}$ except near the edge of the TC, as found in previous studies (buffett96a; buffett98) when the radial field strength at the ICB is above a few mT.
+Inside the TC, $\Omega_{f}$ is locked with $\Omega_{i}$ except near the edge of the TC, as found in previous studies [7, 8].
+The rotation of the mantle, $\Omega_{m}$, is opposite that of the inner core, and its amplitude matches that of the MICG mode under the assumption that the whole of the TC oscillates as a rigid body, $\Omega_{m}=-(C_{tc}/C_{m})\Omega_{i}$.
+(Note that $\Omega_{m}$ is increased by a factor 100 on Figure 2 so that it reaches a similar scale as $\Omega_{i}$ and $\Omega_{f}$).
 The departure of $\Omega_{f}$ from $\Omega_{i}$ near the edge of the TC implies a slight reduction in the effective moment of inertia of the TC taking part in the MICG mode and explains why its frequency is slightly higher than its theoretical prediction (see Figure 1a).
 Also visible in Figure 2a are the Alfvén waves generated from the TC outward.
-With $\overline{B}_{s}=0.1$ mT, their wave speed is sufficiently slow that they get attenuated before they reach the equator of the CMB ($s=r_{f}$).
+With $\overline{B}_{s}=0.1$ mT, and at the frequency of the MICG mode, their wavelength is sufficiently small that their typical time of attenuation is $\tau_{d}=25$ yr.
+This is shorter than $\tau_{A}$, so $Lu<1$, and Alfvén waves get attenuated before they reach the equator of the CMB ($s=r_{f}$).
+These waves carry angular momentum but their short wavelengths imply a small net angular momentum when integrated over the ROTC, and the latter plays a negligible role in the MICG angular momentum budget.
+Figure 1: Frequency and $Q$ of a subset of the normal modes obtained with $\overline{\Gamma}=3\times 10^{20}$ N m, (a) $\overline{B}_{s}=0.1$, (b) $\overline{B}_{s}=0.15$, and (c) $\overline{B}_{s}=0.3$, and different choices of mantle conductance $G_{m}$ (see legend).
+Black squares show the TO modes in the absence of gravitational coupling: $\overline{\Gamma}=0$.
+In each panel, the predicted $\omega_{micg}$ from Eq.
+(9) is indicated by the pink vertical line.
+The angular velocity structure of the modes identified by a purple outline is shown in Figure 2.
 For comparison, we also show on Figure 1a the modes (black squares) that result when there is no gravitational coupling ($\overline{\Gamma}=0$).
-In this case, the MICG mode is absent from the spectrum of modes; only TO modes are present, and their frequency and $Q$ coincide almost identically with those found with $\overline{\Gamma}\neq 0$.
+In this case, the MICG mode is absent from the spectrum of modes; only TO modes are present, and their frequencies and $Q$ coincide almost identically with those found with $\overline{\Gamma}\neq 0$.
 These results show that when $B_{s}$ within the core is sufficiently weak, the MICG mode emerges as an independent mode of the system, along with a set of TO modes, and the latter are only weakly influenced by gravitational coupling.
-Let us now increase $B_{s}$ while keeping $\overline{\Gamma}=3\times 10^{20}$ N m and $G_{m}=10^{8}$ S.
+For a mantle conductance of $G_{m}=10^{8}$, the typical EM attenuation time of Alfvén waves is $\tau_{em}=119$ yr.
+We also show on Figure 1a the set of modes for three other choices of $G_{m}$: $2\times 10^{8}$ S ($\tau_{em}=60$ yr), $4\times 10^{8}$ S ($\tau_{em}=30$ yr) and $6\times 10^{8}$ S ($\tau_{em}=20$ yr).
+The $Q$ of all modes is reduced because of the enhanced attenuation from EM coupling at the CMB.
+In each case, the MICG mode is present as an independent mode because $Lu<1$, while $Lu_{em}$ is also <1<1 for the 3 cases with higher mantle conductance.
+Let us now increase $B_{s}$ while keeping $\overline{\Gamma}=3\times 10^{20}$ N m and restoring $G_{m}=10^{8}$ S.
 We compute the modes for $\overline{B}_{s}=0.15$ mT (case $b$) and $\overline{B}_{s}=0.3$ mT (case $c$).
 The results are shown on Figure 1 (panels b and c, respectively), where we again show the comparison with the modes computed with $\overline{\Gamma}=0$.
-As $\overline{B}_{s}$ is increased, the MICG mode is no longer a distinct, independent mode of the system.
-Instead, it merges into the spectrum of TO modes, although the frequency and $Q$ of these TO modes is altered by gravitational coupling.
-To understand why, Figures 2b and 2c show the profiles of $\Omega_{f}$ for the modes close to the MICG frequency identified by a purple outline in panels $b$ and $c$ of Figure 1.
-With the larger $B_{s}$-field, Alfvén waves generated at the TC travel faster and can propagate further before they get attenuated.
-While for case $b$ Alfvén waves are just barely able to make it through the entire width of the ROTC, they easily reach $s=r_{f}$ in case $c$.
+Figures 2b and 2c show the profiles of $\Omega_{f}$ for the modes close to the MICG frequency identified by a purple outline in panels $b$ and $c$ of Figure 1.
+With a larger $B_{s}$-field, Alfvén waves generated at the TC travel faster, their wavelengths at frequencies close to $\omega_{micg}$ is longer (increasing their ohmic diffusion time $\tau_{d}$), so they propagate further before they get attenuated.
+The travel and ohmic diffusion times for case $b$ are $\tau_{A}=56$ yr, $\tau_{d}=58$ yr and those for case $c$ are $\tau_{A}=28$ yr, $\tau_{d}=235$ yr.
+In each case, $Lu>1$ (though barely for case $b$) and (with $\tau_{em}=119$ yr) $Lu_{em}>1$; Alfvén waves can travel the width of ROTC before they get attenuated.
 Consequently, the structure of Alfvén waves within the whole of the ROTC must be compatible with their end-condition at $s=r_{f}$ and they thus have a feedback on the oscillating motion of the TC.
-In other words, the oscillating TC is now taking part in TO modes, although these are modified by gravitational coupling.
-Figure 1: Frequency and $Q$ of a subset of the normal modes obtained with (a) $\overline{B}_{s}=0.1$, (b) $\overline{B}_{s}=0.15$, and (c) $\overline{B}_{s}=0.3$ and different choices of coupling parameters (Pale blue circles: $\overline{\Gamma}=3\times 10^{20}$ N m, $G_{m}=10^{8}$ S; black squares: $\overline{\Gamma}=0$ N m, $G_{m}=10^{8}$; red diamonds: $\overline{\Gamma}=3\times 10^{20}$ N m, $G_{m}=10^{9}$ S).
-In each panel, the predicted $\omega_{micg}$ from Eq.
-(9) is indicated by the vertical orange line.
-The angular velocity structure of the modes identified by a purple outline is shown in Figure 2.
-The attenuation of Alfvén waves can be achieved by reducing their speed, as illustrated by case $a$, or by increasing their attenuation rate from EM coupling at the CMB.
-To illustrate this, we carried out a set of experiments where we have kept $\overline{\Gamma}=3\times 10^{20}$ N m, the same three values of $\overline{B}_{s}$ as for cases $a$, $b$ and $c$ above, but used an increased mantle conductance of $G_{m}=10^{9}$ S.
-The modes that result from these experiments are shown on Figure 1 as red diamonds.
-The $Q$ of all modes is reduced because of the enhanced attenuation caused by the stronger EM coupling at the CMB.
-In each case though, a distinct MICG mode remains present in the mode spectrum, even with $\overline{B}_{s}=0.3$ mT.
-The above analysis allows us to refine the criteria that determines whether the MICG remains an independent mode of the system.
-The condition is whether Alfvén waves launched by the oscillating TC get attenuated before they can reach $s=r_{f}$.
-When Alfvén waves are attenuated, they do not have a feedback on the motion of the TC, and their oscillating nature imply a weak net angular momentum change of the ROTC.
-In this sense, the ROTC plays a passive role, and the MICG mode, for which the angular momentum exchange remains dominantly between the mantle and the TC via gravitational coupling, is a possible independent mode of oscillation.
-In contrast, if Alfvén waves launched at the TC can travel through the entire width of the ROTC, the motion of the TC must be compatible with the allowed structure of TO modes.
-Figure 2: The angular velocity of the inner core $\Omega_{i}$ (thick orange) and of the fluid core $\Omega_{f}$ (blue) as a function of cylindrical radius ($s/r_{f}$) for the modes identified by a purple outline on the respective panels of Figure 1.
+In each case, the MICG mode is no longer a distinct, independent mode of the system.
+Instead, it merges into the spectrum of TO modes.
+Figures 1b and 1c also show the resulting modes when the mantle conductance is increased to $2\times 10^{8}$ S ($\tau_{em}=60$ yr), $4\times 10^{8}$ S ($\tau_{em}=30$ yr) and $6\times 10^{8}$ S ($\tau_{em}=20$ yr).
+By progressively decreasing $\tau_{em}$, a point is reached when Alfvén waves get attenuated before they can travel the width of the ROTC (that is, $Lu_{em}$ becomes smaller than 1), and a distinct MICG mode re-emerges in the spectrum of modes.
+With $\overline{B}_{s}=0.15$ mT ($\tau_{A}=56$ yr), the MICG is almost fully out of the TO branch for $G_{m}=2\times 10^{8}$ S ($Lu_{em}\sim 1$), and clearly separated for $G_{m}>4\times 10^{8}$ S .
+With $\overline{B}_{s}=0.3$ mT ($\tau_{A}=28$ yr), a larger $G_{m}=6\times 10^{8}$ S is required for $Lu_{em}<1$ and for the MICG mode to split from the TO branch.
+Figure 2: The angular velocity of the inner core $\Omega_{i}$ (thick light blue), mantle $\Omega_{m}\times 100$ (thick orange) and of the fluid core $\Omega_{f}$ (thin dark blue) as a function of cylindrical radius ($s/r_{f}$) for the modes identified by a purple outline on the respective panels of Figure 1.
+The dashed black line corresponds to the prediction $\Omega_{m}=-(C_{tc}/C_{m})\Omega_{i}$ (multiplied by 100) when the exchange of angular momentum is between the mantle and whole of TC.
 In each panel, the phase is chosen when $\Omega_{i}$ is maximum in the prograde direction and the amplitude is normalized such that $\max|\Omega_{f}|=1$.
-Since we expect the $B_{s}$-field strength in the core to be of the order of a few mT, Alfvén waves should propagate through the ROTC in a short timescale of a couple of years, more rapidly than their timescale of attenuation by EM coupling at the CMB (which should be of the order of a few decades e.g.
-dumberry08c; gillet10).
-The MICG mode should then not exist as a distinct mode of oscillation; rather, we expect it to be absorbed into the spectrum of TO modes, although the TO modes are influenced by gravitational coupling.
-To further explore how gravitational coupling affects TO modes, let us restore $G_{m}=10^{8}$ S and now set $\overline{B}_{s}=3$ mT.
-With these choices, the MICG mode is not an independent mode and the first TO harmonic has a period of approximately 6 years.
-Figure 3 shows how the period (panel a) and $Q$ (panel b) of the first four TO harmonics are altered as a function of $\overline{\Gamma}$.
-Figure 4 shows how the structure of the first three TO harmonics differ in a weak ($\overline{\Gamma}=10^{19}$ N m) versus a strong ($\overline{\Gamma}=10^{22}$ N m) gravitational coupling regime.
-The oscillation of the TC – entrained by the TO modes – is only weakly resisted by gravitational coupling when the latter is weak.
-In contrast, when gravitational coupling is strong, the oscillation of the TC is severely restricted.
-Using standing waves on a string as an analogy, a strong gravitational coupling changes the boundary condition on $\Omega_{f}$ from a free-end at $s\rightarrow 0$ to one more akin to a fixed-end at $s<s_{tc}$.
-Because the latter implies a shorter mode wavelength, the period of the modes is shorter, as seen in Figure 3a.
-Figure 3a also shows that the point of transition of every TO mode between the weak and strong gravitational coupling regime is set by the period of the MICG mode.
-In other words, the TC remains gravitationally quasi-locked to the mantle for a given TO mode if its period is much longer than the MICG period.
-Hence, even though the MICG mode is no longer an individual mode of the system, its period retains an influence in the characteristics of the resulting TO modes.
-When the period of a TO mode approaches the MICG period, a resonance effect occurs, leading to a higher $Q$ (Figure 3b).
-The degree of resonant amplification depends on the attenuation of the TO modes caused by EM damping at the CMB but also by the viscous relaxation time of the inner core $\tau$.
-When the latter is of the same order or shorter than the period of a given TO mode, $Q$ is reduced.
-This is shown in Figure 3b, where we have computed $Q$ for $\tau=100$, $10$ and $3$ years.
+The above analysis shows that the MICG only remains an independent mode when Alfvén waves launched by the oscillating TC get attenuated before they reach $s=r_{f}$ (either because $Lu<1$ or $Lu_{em}<1$).
+If so, they do not have a feedback on the motion of the TC, and the MICG mode can exist independently from the TO modes.
+In contrast, when Alfvén waves can travel through the entire width of the ROTC, the motion of the TC must be compatible with the structure of TO modes, and the MICG mode gets absorbed in the spectrum of TO modes.
+Alfvén waves may be partially or fully absorbed once they reach the equator of the CMB [44, 24], but the key factor is whether they can reach this point.
+The $B_{s}$-field chosen for cases $b$ and $c$ above remains weak (smaller than 0.3 mT).
+The TO harmonics at periods close to the MICG mode have short wavelengths and the net angular momentum of the ROTC for these modes remains small.
+The angular momentum budget for these TO modes remains primarily an exchange between the mantle and the whole of the TC via gravitational coupling.
+Figures 2b and 2c show that $\Omega_{m}$ indeed remains approximately equal to $(C_{tc}/C_{m})\Omega_{i}$ for these selected modes.
+Hence, these TO modes share the same angular momentum dynamics as the MICG mode.
+Although presented differently, our results are in agreement with those of [37].
+In their study, which uses a similar model, they do not compute modes directly.
+Rather, they impose a forcing on the inner core, and monitor the response of the mantle.
+The frequencies at which the angular motion of the mantle is amplified mark the periods of the normal modes.
+With a uniform $B_{s}$-field of 0.4 mT (similar to our case $c$), and with no EM coupling at the CMB, they find (as we do) a set of TO modes amplified in the broad vicinity of the MICG frequency (see their Figure 4).
+This amplification leaves a broad peak in the LOD at the period of the MICG mode [39].
+We now explore how the MICG mode affects the TO modes in a strong $B_{s}$-field regime.
+We restore $G_{m}=10^{8}$ S and set $\overline{B}_{s}=3.3$ mT, typical of the field strength expected inside the core [25].
+With these choices, the fundamental TO mode has a period of approximately 6 years.
+Alfvén waves can readily propagate across the entire ROTC in a short timescale of $\tau_{A}=2.5$ yr.
+This is much faster than $\tau_{em}$ and also much faster than $\tau_{d}$ which (for a period of 6 yr) is now close to $3\times 10^{4}$ yr.
+Hence, both $Lu\gg 1$ and $Lu_{em}\gg 1$, and, accordingly, the spectrum of modes should not include an independent MICG mode; we have verified that this is indeed the case for any choices of $\overline{\Gamma}$.
+With our chosen mantle conductance, $Lu_{m}=0.34$; this is sufficiently high that Alfvén waves are partly absorbed at the equator of the CMB, yet low enough that distinct spectral peaks at the periods of the TO modes would remain visible in the LOD spectrum, compatible with observations [24].
 Figure 3: The period (a) and $Q$ (b) of the first four TO harmonics (coloured lines, legend) as a function of $\overline{\Gamma}$.
@@
 In (b), results are shown for $\tau=100$ years (solid lines), $\tau=10$ years (thin solid lines) and $\tau=3$ years (dashed lines).
-Figure 4: The angular velocity as a function of cylindrical radius ($s/r_{f}$) of the (a) first, (b) second, and (c) third TO harmonics, for $\overline{\Gamma}=10^{19}$ N m (blue lines for $\Omega_{f}$, thick light blue lines for $\Omega_{i}$) and $\overline{\Gamma}=10^{22}$ N m (red lines for $\Omega_{f}$, thick orange lines for $\Omega_{i}$).
-In each case, the phase is chosen when $\Omega_{i}$ is maximum in the prograde direction and the amplitude is normalized such that $\max|\Omega_{f}|=1$.
+Figure 4: The angular velocity as a function of cylindrical radius ($s/r_{f}$) of the (a, d) first, (b, e) second, and (c, f) third TO harmonics for $\overline{\Gamma}=10^{19}$ N m (left column) and $\overline{\Gamma}=10^{22}$ N m (right column).
+Thin blue lines: $\Omega_{f}$, thick light blue lines: $\Omega_{i}$ and thick orange lines: $\Omega_{m}\times 100$.
+The dashed black line corresponds to the prediction $\Omega_{m}=-(C_{tc}/C_{m})\Omega_{i}$ (multiplied by 100).
+The dashed red line corresponds to the prediction of $\Omega_{m}=-L_{rotc}/C_{m}$ (multiplied by 100), where $L_{rotc}$ is the net angular momentum of the ROTC.
+In each panel, the phase is chosen when $\Omega_{i}$ is maximum in the prograde direction and the amplitude is normalized such that $\max|\Omega_{f}|=1$.
+Figure 3 shows how the periods (panel a) and $Q$ (panel b) of the first four TO harmonics are altered as a function of $\overline{\Gamma}$.
+Figure 4 shows how the structure of the first three TO harmonics differ in a weak ($\overline{\Gamma}=10^{19}$ N m) versus a strong ($\overline{\Gamma}=10^{22}$ N m) gravitational coupling regime.
+When gravitational coupling is weak ($\overline{\Gamma}=10^{19}$ N m), the oscillation of the TC – entrained by the TO modes – is only weakly resisted by gravitational coupling.
+We also show on Figure 4 the predictions of $\Omega_{m}$ based on an angular momentum exchange with the TC ( $\Omega_{m}=-(C_{tc}/C_{m})\Omega_{i}$) and with the ROTC ($\Omega_{m}=-L_{rotc}/C_{m}$, where $L_{rotc}$ is the net angular momentum of the ROTC).
+The sum of these predictions matches $\Omega_{m}$ well, reflecting that the core-mantle angular momentum is conserved.
+But neither provides a good fit to $\Omega_{m}$, which has a much smaller amplitude.
+This indicates that the angular momentum budget for these modes is primarily an exchange between the TC and the ROTC; $\Omega_{m}$ reflects the small imbalance in the angular momenta of the TC and the ROTC.
+When gravitational coupling is strong ($\overline{\Gamma}=10^{22}$ N m), the oscillation of the TC is more restricted.
+Using standing waves on a string as an analogy, a strong gravitational coupling changes the boundary condition on $\Omega_{f}$ from a free-end at $s\rightarrow 0$ to one more akin to a fixed-end at $s<s_{tc}$.
+Because the latter implies a shorter mode wavelength, the period of the TO modes are shorter (see Figure 3a).
+The angular momentum budget is now dominantly an exchange between the mantle and the ROTC (although the TC plays a non-negligible role).
+Whether $\overline{\Gamma}$ is weak or strong, these TO modes have a different angular momentum dynamics than the MICG mode.
+Instead, the ROTC is involved in the angular momentum budget and it is the whole of the core that exchanges its angular momentum with the mantle.
+Figure 3a also shows that the point of transition between the weak and strong gravitational coupling regime for every TO mode is set by the period of the MICG mode.
+Hence, even though the MICG mode is no longer an individual mode of the system, its period retains an influence in the characteristics of the resulting TO modes.
+When the period of a TO mode approaches the MICG period, a resonance effect occurs, leading to a higher $Q$ (Figure 3b).
+Note that the level of amplification is larger when the MICG period matches higher TO harmonics, consistent with results obtained in the weak $B_{s}$-field regime.
+When the MICG period is longer than the period of the fundamental TO harmonic, the level of amplification is limited.
+The amplification also depends on the attenuation of the TO modes caused by EM damping at the CMB and by the viscous relaxation time of the inner core $\tau$.
+When the latter is of the same order or shorter than the period of a given TO mode, $Q$ is significantly reduced.
+This is shown in Figure 3b, where we have computed $Q$ for $\tau=100$, $10$ and $3$ years.
 5 Discussion and conclusions
-As our results show, when the magnetic field strength in the bulk of the outer core is of the order of a few mT as predicted from dynamo models (christensen06; aubert09), the TC is strongly coupled by EM stress to the fluid cylinders in the ROTC.
-Motion in the latter take the form of Alfvén waves.
-If Alfvén waves launched at the TC can traverse through the entire width of the ROTC before being attenuated, the MICG mode is no longer a distinct independent mode of the coupled Earth system.
-Instead, it gets absorbed into the spectrum of normal modes of TO.
-The MICG period retains a meaning, acting as a point of resonance for TO modes, and marking the transition from a TO mode in which the oscillating motion of the TC is weakly impeded by gravitational coupling to one where the TC remains quasi-locked to the mantle.
-Though presented differently, our results are in agreement with those of mound03.
-In their study, which uses a broadly similar model as ours, they do not compute the modes of oscillations directly.
-Rather, they impose a forcing on the inner core, and monitor the response of the mantle angular velocity; frequencies at which the motion of the latter is amplified mark the location of normal modes.
-With a uniform $B_{s}$-field of 0.4 mT (similar to our case $c$), and with no EM coupling at the CMB, they find a set of resonances which are amplified in the broad vicinity of the MICG frequency associated with TO modes (see their Figure 4).
-This is in agreement with our results: we also find that TO modes have a higher $Q$ when their periods is in the vicinity of the MICG period.
-mound03 interpreted these results to imply that the MICG mode is complicated by its interaction with TO modes.
-By directly computing the modes of the coupled system, we have refined this interpretation; the MICG mode is no longer a distinct independent mode, it gets integrated into the spectrum of TO modes when Alfvén waves are weakly attenuated.
-With a $B_{s}$ strength in the core expected to be of the order of a few mT, Alfvén waves can traverse the core in only a few years (gillet10).
-For the MICG mode to remain an independent mode of the coupled Earth system, the attenuation time of Alfvén waves by coupling at the CMB (EM or otherwise) would then need to be of the order of one year or less.
-As an indication, such a fast attenuation would require a mantle conductance larger than $10^{9}$ S, so large in fact that the gravest TO modes and the MICG mode would all have a $Q$ below 1.
-This conclusion is important regarding the interpretation of the 5.9 year oscillation in the LOD, which includes a clear contribution from core (e.g.
-pfeffer23; cazenave25).
-A number of studies have suggested that this signal may represent the signature of the MICG mode (mound06; chao17; ding18; duan18; shih21).
-Our results show that this interpretation, which involves the ROTC playing a passive role, cannot be correct when the magnetic field strength inside the core is a few mT.
-Consequently, a measure of the gravitational coupling constant $\overline{\Gamma}$ cannot be inferred on the basis of the 5.9 year LOD signal, and in turn, it cannot be used as an indirect constraint on the amplitude of the degree 2 order 2 mantle mass anomalies.
-Interannual LOD changes involving the core must then be the manifestation of either TO modes, or any other free or driven core flow motion.
-This is consistent with the results of gillet10 who showed that, for the 5.9 year LOD signal, the mantle exchanges its angular momentum with the fluid core, and dominantly the ROTC given that latter has a much larger moment of inertia than the TC.
-The most recent estimates of the gravitational coupling constant $\overline{\Gamma}$ based on mantle flow models suggest a range of $0.3-2\times 10^{20}$ N m (davies14), corresponding to a range of MICG periods of 7 to 19 years.
+Our results show that if Alfvén waves launched at the TC can travel across the entire width of the ROTC before being attenuated, the MICG mode is no longer a distinct independent mode of the core-mantle system.
+For a magnetic field strength in the bulk of the outer core equal to a few mT, as predicted by dynamo models [14, 3], Alfvén waves can readily traverse the core in only a few years [25], much shorter than their attenuation time.
+The MICG mode should be absorbed into the spectrum of TO modes.
+The MICG period retains a dynamical influence, as it acts as a point of resonance for TO modes, and marks the transition from a TO mode in which the oscillating motion of the TC is weakly affected by gravitational coupling to one in which the motion of the TC is strongly impeded by it.
+For the MICG mode to remain an independent mode of the core-mantle system, the attenuation time of Alfvén waves by coupling at the CMB (EM or otherwise) would need to be of the order of one year or less.
+As an indication, such a fast attenuation requires a mantle conductance larger than $10^{9}$ S, so large in fact that the gravest TO modes and the MICG mode would all have a $Q$ below 1.
+Hence, the very fact that TO modes are excited to an observable level points to a limit in the mantle conductance which is well below $10^{9}$ S [24].
+In addition, such a large conductance would introduce a significant delay between the time a magnetic signal is generated at the CMB and its observation at the surface [33, e.g.], yet no evidence of such a delay is observed [30, 26, e.g.].
+The latest results from mineral physics and seismology also point to an upper bound on the mantle conductance well below $10^{9}$ S.
+The electrical conductivity of FeO at CMB conditions is in the range of $10^{4}-10^{5}$ S m-1 [41, 29], and while a thin layer of approximately 2 km of such iron-enriched material is supported by normal modes [43], a thicker layer of a few tens of km – such as would be required for a conductance $10^{9}$ S – is not.
+Given the limits on the possible mantle conductance, it is difficult to escape the conclusion that the MICG mode should be absorbed in the spectrum of TO modes.
+This conclusion is important regarding the interpretation of the 5.9 year oscillation in the LOD, which includes a clear contribution from the core [42, 11, e.g.].
+A number of studies have suggested that this signal may represent the signature of the MICG mode [39, 12, 17, 18, 45].
+This interpretation rests on the assumption that the magnetic field in the core is weak (equal to a fraction of a mT).
+In such a case, the resonant excitation of a cluster of high TO harmonics with periods close to that of the MICG produces a distinct LOD peak at the MICG period [37, 39].
+The angular momentum budget for these TO modes is primarily an exchange between the mantle and the whole of the TC, just as it is for the MICG mode.
+However, when the magnetic field strength inside the core is a few mT (as we expect), the subdecadal period range is now populated by the first few TO harmonics.
+These have distinct, separated periods and their angular momentum dynamics feature a leading order contribution from the ROTC; they do not behave like the MICG mode.
+These TO modes have a higher $Q$ if their periods are close to the MICG period, but the peaks in the LOD that they may leave correspond to the distinct periods of the TO modes, not the period of the MICG.
+Hence, for a strong magnetic field inside the core of a few mT, the 5.9 year LOD signal cannot represent a direct signature of the MICG mode.
+Instead, it must be a signature of the fundamental TO mode, or any other free or driven core flow motion.
+This is consistent with the results of [25] who showed that, for the 5.9 year LOD signal, the mantle exchanges its angular momentum with the whole of the core, not the TC alone.
+The most recent estimates of the gravitational coupling constant $\overline{\Gamma}$ based on mantle flow models suggest a range of $0.3-2\times 10^{20}$ N m [15], corresponding to a range of MICG periods between 7 and 19 years.
 Assuming a $B_{s}$-field of approximately 3 mT or larger, all TO modes should then have a period shorter than the MICG period.
 This implies that TO modes should be weakly influenced by gravitational coupling in the sense that oscillations of the TC region should not be severely restricted by the mantle.
-Furthermore, the period of the TO modes, especially those of the first few harmonics, may be sufficiently close to the MICG period that they may be resonantly amplified.
-We thus expect that, at subdecadal periods, the TC (including the inner core) can undergo oscillating motion with respect to the mantle at a similar angular velocity amplitude as cylinders outside the TC, whether driven by TO modes or by any other core flow.
-At a period of 6 years, and based on the observed fluctuations of the magnetic field, zonal flow fluctuations inside the TC have an amplitude of approximately $v_{\phi}\sim 0.4$ km yr-1 (gillet10).
-Using this as a representative measure, this corresponds to longitudinal angular fluctuations of the inner core of approximately $0.02^{\circ}$ (lecomte23).
-Our results also have important implications for the interpretation of the subdecadal fluctuations in inner core rotations recently inferred by seismic observations, notably in the study of wang22 which suggests a super-rotation of the inner core of $\sim 0.29^{\circ}$ between 1971 and 1974 and a sub-rotation of $\sim 0.1^{\circ}$ between 1969 and 1971.
-wang22 further showed that the amplitude and phase of these inferred inner core rotation changes ($\Omega_{i}$) are consistent with those reconstructed from the 5.9 year LOD signal under the assumption that the latter is caused by the MICG mode.
-We first note that the inferred amplitude of inner core oscillations, if part of the MICG mode, would then imply similar fluctuations in zonal core flows inside the TC, and these are much larger than the approximate upper bound of $0.02^{\circ}$ stated above.
-It is possible that zonal core flows are not sufficiently well resolved inside the TC, so their amplitude may be larger.
-Alternatively, a part of the seismic signal could be explained by changes in the ICB topography (wen06; yao19), including from viscous relaxation (vidale25).
-In any case, its large amplitude notwithstanding, our results indicate that subdecadal changes in $\Omega_{i}$ cannot be associated with the MICG mode, contrary to the interpretation made by wang22.
-Instead, we favour an interpretation of subdecadal $\Omega_{i}$ changes entrained by TO modes or, more generally, by Alfvén waves.
+We thus expect that, at subdecadal periods, the TC (including the inner core) can undergo oscillating motion with respect to the mantle at a similar angular velocity amplitude as cylinders outside the TC.
+At a period of 6 years, and based on the observed fluctuations of the magnetic field, zonal flow fluctuations inside the TC have an amplitude of approximately $v_{\phi}\sim 0.4$ km yr-1 [25].
+Using this as a representative measure, this corresponds to longitudinal angular fluctuations of the inner core of approximately $0.02^{\circ}$ [34].
+Our results also have important implications for the interpretation of the subdecadal fluctuations in inner core rotations recently inferred by seismic observations, notably in the study of [48] which suggests a super-rotation of the inner core of $\sim 0.29^{\circ}$ between 1971 and 1974 and a sub-rotation of $\sim 0.1^{\circ}$ between 1969 and 1971.
+[48] further showed that the amplitude and phase of these inferred inner core rotation changes ($\Omega_{i}$) are consistent with those reconstructed from the 5.9 year LOD signal under the assumption that the latter is caused by the MICG mode.
+The authors have themselves challenged their own interpretation in a subsequent study focused on the time window 1991-2023, in which a 6-yr inner core oscillation of a similar magnitude is not observed [49].
+Based on our results, we favour an interpretation of subdecadal changes in $\Omega_{i}$ entrained by TO modes or, more generally, by Alfvén waves.
 In this case, unlike for the MICG mode, a prediction of $\Omega_{i}$ cannot be constructed directly from the LOD signal but instead depends on many factors, including the relative strengths of the EM and gravitational torque on the mantle and the structure of Alfvén waves.
 It is of course possible that the phase of $\Omega_{i}$ entrained by Alfvén waves is similar to that predicted from the LOD signal on the basis of the MICG mode.
-Finally, we have noted that the MICG period can act as a resonance to amplify TO modes.
-More generally any driven motion of the TC at a period that approaches that of the MICG would be amplified.
-If $\overline{\Gamma}\approx 3\times 10^{19}$ N m, the MICG period is approximately 20 years.
-If the inner core viscous relaxation time is of the same order or longer than this period, some of the observed multidecadal LOD changes caused by the core (e.g.
-finlay23) may then result from an amplification of driven zonal core flow motion (e.g.
-more18) due to their close temporal proximity to the MICG period.
+However, the amplitude of the subdecadal changes $\Omega_{i}$ should be limited to approximately $0.02^{\circ}$, as stated above, much smaller than the amplitudes inferred by [48].
+Although our focus here has been on the MICG mode, our model may be adapted to monitor the angular momentum response of the core-mantle system to a specified forcing in the core, either at a given frequency [37, 38, e.g.] or for a stochastic forcing [24, e.g.].
+The parameter space of $\overline{\Gamma}$, $\tau$ and $G_{m}$ may be surveyed to determine the range of values that allow to best reproduce the characteristics of the observed Alfvén waves, while also matching the amplitude and phase of the LOD changes at subdecadal and decadal timescales.
+This would provide constraints on these parameters, which in turn would provide further constraints on the structure, composition and dynamics of the lower mantle and inner core.
 Acknowledgements.
-This work was financially supported by a Discovery Grant from NSERC/CRSNG (Canada). Figures were produced using the GMT software (gmt).
+I wish to thank Jon Mound and an anonymous reviewer for comments and suggestions that have greatly improved the quality of this paper.
+This work was financially supported by a Discovery Grant from NSERC/CRSNG (Canada).
+Figures were produced using the GMT software [50].
 Data availability
-The codes necessary to reproduce all of our numerical experiments are freely accessible at the following data repository: dumberry25data.
+Codes to reproduce all numerical experiments are freely accessible at the following data repository: [19].
 Competing interests
 The author has no competing interests.
-[…]